### v2

# arXiv:2601.06128v2[math.GM] 01 Mar 2026

Abstract.
Fix $\Lambda>0$ and consider trace-normed $2\times 2$ canonical systems $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}Y^{\prime}(s)=zH(s)Y(s)$ on $0,\Lambda]$, extended by the free tail $H(s)=\tfrac{1}{2}I$ for $s\geq\Lambda$. The Weyl–Titchmarsh coefficient $m_{H,\Lambda}$ and its Cayley transform $v$ encode the spectral measure. We study the finite sampling operator (the seam map) that sends $H$ (or a finite–dimensional parameter $\theta$) to the values $v(z_{k})$ at points $z_{k}=x_{k}+i\eta\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, $k=1,\dots,M$. For piecewise–constant (block) Hamiltonians, with block length $\ell=\Lambda/N$, we analyze stability and optimal sampling.
We prove three types of results. (1) The free tail forces the transfer matrix at $\Lambda$ to have entire entries of exponential type $\leq\Lambda$; consequently $m_{H,\Lambda}$ and $v$ extend meromorphically to $\mathbb{C}$, showing that the free-tail class is a strict spectral subclass. (2) Fixed–height data is intrinsically smoothed: $\operatorname{Im}m(\cdot+i\eta)$ is the Poisson extension of the spectral measure, yielding an exponential minimax obstruction to recovering oscillatory spectral features from finitely many noisy samples. (3) Linearization at the free Hamiltonian identifies seam sampling with an evaluation operator in a Paley–Wiener/de Branges space, so stability is governed by kernel–frame bounds; shifted equispaced designs give optimal tight frames and explicit conditioning estimates. Finally, a perturbation formula expresses $Dm$ as an $L^{2}_{H}$ pairing with $\Delta H$, linking depth attenuation to decay of Weyl solutions.
Key words and phrases:
canonical systems, Weyl–Titchmarsh function, Schur transform, inverse problems, minimax bounds, sampling design, instability, de Branges spaces, Carleson measures, prolate spheroidal functions
2020 Mathematics Subject Classification:
34A55, 34B20, 34L05, 47A10, 30H10
1. Introduction and main results
1.1. Setup
Let $\Lambda>0$ and consider the $2\times 2$ canonical system
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,Y^{\prime}(s)=z\,H(s)\,Y(s),\qquad s\in[0,\Lambda],$ (1)
where $H(s)\in\mathbb{R}^{2\times 2}_{\mathrm{sym}}$ is measurable, $H(s)\succeq 0$ a.e., and $\operatorname{tr}H(s)=1$ a.e. We attach the constant tail
$H(s)\equiv\tfrac{1}{2}I,\qquad s\geq\Lambda,$ (2)
and denote by $m_{H,\Lambda}(z)$ the corresponding Weyl coefficient, normalized by $Y(0)=(1,m)^{\top}$. Then $m_{H,\Lambda}$ is Herglotz–Nevanlinna (analytic on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ with $\operatorname{Im}m>0$); see, e.g., [9] (up to standard sign and normalization conventions). Its Cayley transform
$v_{H,\Lambda}(z):=\frac{m_{H,\Lambda}(z)-i}{m_{H,\Lambda}(z)+i}$ (3)
is a Schur function on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (equivalently, after a Möbius map from $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ to the unit disk); see, e.g., [10].
Remark 1.1 (Complex notation for real-symmetric traceless directions).
We often write real-symmetric traceless perturbations using a complex parameter $q=a+ib\in\mathbb{C}\simeq\mathbb{R}^{2}$ via
$\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}=\begin{pmatrix}a&b\\ b&-a\end{pmatrix}$.
This is purely notational: all Hamiltonians and perturbations remain real-symmetric.
Fixing a height $\eta>0$ and real nodes $x_{1},\dots,x_{M}\in\mathbb{R}$, we set $z_{k}:=x_{k}+i\eta$. Sampling defines the seam map
$\mathcal{S}(\theta):=\bigl(v_{H(\theta),\Lambda}(z_{k})\bigr)_{k=1}^{M}\in\mathbb{D}^{M},$ (4)
where $\theta$ parametrizes a finite-dimensional Hamiltonian family.
In words, $\mathcal{S}(\theta)$ records the values of the Schur transform of the Weyl function at height $\operatorname{Im}z=\eta$ and at real sampling locations $x_{1},\dots,x_{M}$.
Fixed-height data is spectrally smoothed: $\operatorname{Im}m(\cdot+i\eta)$ is a Poisson smoothing of the spectral measure (Corollary˜10.9), yielding an exponential de-smoothing barrier in frequency (Theorem˜10.12). At the free base point $H_{0}\equiv\tfrac{1}{2}I$, the Fréchet derivative of $v$ is an explicitly weighted Fourier–Laplace transform (Theorem˜3.5), exhibiting exponential suppression of sensitivity to perturbations supported deeper in the interval. These two mechanisms are logically independent: the first is global and base-point-free, while the second is a local (linearized) depth effect that sharpens to a design-minimax barrier in finite-dimensional models. Here “depth” refers to the $s$-location of a perturbation: contributions supported closer to $s=\Lambda$ are “deeper” in the interval.
1.2. Main results
We treat fixed-height values of the Weyl–Schur function as a spectral sampling problem for canonical systems. The main results can be grouped as follows.
  1. (i)
    Free-tail canonical systems: a strict class and a transfer-matrix characterization. If $H$ is trace-normed on $0,\Lambda]$ and equals $\tfrac{1}{2}I$ for $s\geq\Lambda$, then the Weyl function is a quotient of entire functions and hence extends meromorphically to $\mathbb{C}$ (Theorem˜10.2). More intrinsically, we give an if-and-only-if characterization of this free-tail Weyl class in terms of entire $J$-contractive transfer matrices of finite exponential type (Theorem˜10.6). We also show strictness by exhibiting a Herglotz function with unavoidable logarithmic branching that cannot arise from any free-tail system (Remark˜10.3).
  2. (ii)
    A base-point-free exponential obstruction from spectral smoothing. Independently of any reference Hamiltonian, fixing $\operatorname{Im}z=\eta$ smooths spectral data: $\operatorname{Im}m(\cdot+i\eta)$ is the Poisson extension of the spectral measure (Corollary˜10.9). From this we prove a deterministic minimax lower bound showing that recovering oscillatory spectral features from finitely many noisy fixed-height samples is exponentially ill-conditioned in the feature frequency (Theorem˜10.12).
  3. (iii)
    Linearized seam sampling as evaluation in Paley–Wiener/de Branges spaces. At the free Hamiltonian $H_{0}\equiv\tfrac{1}{2}I$, the Fréchet derivative $Dv_{H_{0},\Lambda}$ is a weighted Fourier–Laplace transform (Theorem˜3.5) whose range is a Paley–Wiener/de Branges space (Proposition˜9.1). Sampling is therefore an RKHS evaluation operator, and conditioning is governed by kernel-frame/Gram-matrix bounds (Proposition˜9.3 and Section˜9).
  4. (iv)
    General finite-dimensional tangent models: a basis-dependent depth barrier. For arbitrary finite-dimensional tangent parametrizations $q_{\theta}=\sum_{j=1}^{N}\theta_{j}\,\varphi_{j}$ (not necessarily block-constant), we derive an explicit Jacobian for seam sampling and obtain a universal depth-barrier upper bound on its smallest singular value (Theorem˜4.5).
  5. (v)
    A sharp solvable case (block model): an explicit depth barrier and optimal kernel frames. For block-constant Hamiltonians we factor the Jacobian into a row factor, a Fourier sampling block, and exponential depth weights. This yields universal depth-barrier upper bounds and explicit near-optimal shifted equispaced designs (tight kernel frames) with computable constants (Theorem˜1.3 and Section˜6).
  6. (vi)
    Quantitative local inversion and matched local minimax bounds. In the square sampling regime we prove an explicit inverse function theorem with computable constants and a fixed-point reconstruction scheme, and we match these stability guarantees with deterministic local minimax lower bounds under adversarial noise (Sections˜7 and 8).
  7. (vii)
    A perturbation formula for $m$ as a spectral pairing and an operator corollary. We express $Dm(z)[\Delta H]$ as an $L^{2}_{H}$-pairing against the Weyl solution, making depth attenuation a direct consequence of decay of canonical-system solutions (Section˜2.2). We also record a Dirac-type interpretation: free-tail systems model compactly supported perturbations of the free Dirac operator, and the fixed-height minimax barrier becomes an intrinsic limitation of complex-energy sampling (Section˜11).
While the technical core is an inverse/sampling problem, the main results can be read as intrinsic statements about Weyl/Schur data of canonical systems: (i) imposing a free tail is a genuine spectral restriction, forcing the Weyl/Schur functions to be quotients of entire functions (and hence meromorphic on $\mathbb{C}$), see Remark 10.3; (ii) restriction to a horizontal line $\operatorname{Im}z=\eta$ is a smoothing operator (Poisson extension of the spectral measure), which yields an exponential minimax barrier for any reconstruction from finitely many fixed-height samples, see Theorem 10.12; (iii) in the finite Paley–Wiener/de Branges models underlying our Jacobians, stable sampling is a kernel-frame phenomenon, and tight frames maximize the lower frame bound (Theorems 6.46.15); (iv) quantitative local inversion near the free Hamiltonian follows once these frame bounds are combined with the remainder control developed in Sections 3, 4, and 7.
1.3. Norm conventions
Throughout, for linear maps between Euclidean spaces we use the operator norm induced by $\ell^{2}$: $\left\|A\right\|_{2\to 2}=\sup_{\left\|x\right\|_{2}=1}\left\|Ax\right\|_{2}$.
Matrix-valued $L^{p}$ norms.
For a matrix-valued function $A:(0,\Lambda)\to\mathbb{R}^{2\times 2}$, we write
$\|A\|_{L^{p}(0,\Lambda;\mathrm{op})}:=\begin{cases}\left(\int_{0}^{\Lambda}\|A(s)\|_{\mathrm{op}}^{p}\,ds\right)^{1/p},&1\leq p<\infty,\\[4.0pt] \operatorname*{ess\,sup}_{s\in(0,\Lambda)}\|A(s)\|_{\mathrm{op}},&p=\infty.\end{cases}$
When no ambiguity is possible, we abbreviate $\|A\|_{L^{p}(0,\Lambda)}:=\|A\|_{L^{p}(0,\Lambda;\mathrm{op})}$. For bilinear maps $B:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{m}$ we use
$\left\|B\right\|_{2\to 2\to 2}:=\sup_{\left\|u\right\|_{2}=\left\|v\right\|_{2}=1}\left\|B[u,v]\right\|_{2}.$
For matrices acting on $\mathbb{R}^{2}$ we write $\left\|\cdot\right\|_{\mathrm{op}}$ for the induced operator norm. We write $\sigma_{\min}(A)$ for the least singular value on the domain dimension. Concretely, for a linear map $A:\mathbb{C}^{N}\to\mathbb{C}^{M}$ with $M\geq N$, we set
$\sigma_{\min}(A):=\sigma_{N}(A)=\inf_{\|x\|_{2}=1}\|Ax\|_{2},$
i.e. the smallest singular value among the $N$ (possibly) nonzero singular values. If $\operatorname{rank}(A)<N$ then $\sigma_{\min}(A)=0$.
Remark 1.2 (Constant inflation).
Throughout, we occasionally replace bounds like $e^{|z|\Lambda}$ by $e^{c|z|\Lambda}$ with a larger universal $c$ to simplify downstream expressions; all displayed constants remain explicit and valid.
1.4. Main theorem: sharp depth barrier and near-optimal sampling (block model)
The most explicit design-minimax conclusions occur for block-constant Hamiltonians; see Sections˜5 and 6. Set $\ell:=\Lambda/N$.
Theorem 1.3 (Depth barrier and near-minimax sampling at the free point (block model)).
Fix $\eta>0$ and integers $M\geq N\geq 1$ with $M\geq 2$. Consider the block-constant Hamiltonian model on $0,\Lambda]$ with $N$ blocks of length $\ell=\Lambda/N$, tail-extended by $H\equiv\tfrac{1}{2}I$ on $\Lambda,\infty)$, and let $\mathcal{S}$ be the seam map sampling $z_{k}=x_{k}+i\eta$.
Let $T_{x}:=D\mathcal{S}(0)$ denote the Jacobian at the free Hamiltonian $H_{0}\equiv\tfrac{1}{2}I$.
  1. (i)
    (Universal depth-barrier upper bound) For every design $x\in\mathbb{R}^{M}$,
    $\sigma_{\min}(T_{x})\ \leq\ 2\sqrt{M}\,\cosh\!\Bigl(\frac{\eta\ell}{2}\Bigr)\,e^{-\eta(\Lambda-\ell/2)}.$
  2. (ii)
    (Shifted-DFT lower bound) For the half-shifted equispaced design
    $x_{k}:=\frac{2\pi(k-\tfrac{1}{2})}{M\ell},\qquad k=1,\dots,M,$
    one has
    $\sigma_{\min}(T_{x})\ \geq\ 2\sqrt{M}\,\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2M}\Bigr)}\,e^{-\eta(\Lambda-\ell/2)}.$
  3. (iii)
    (Design-minimax sharpness up to explicit constants) Writing
    $\Sigma^{*}_{M}:=\sup_{x\in\mathbb{R}^{M}}\sigma_{\min}(T_{x}),$
    $\Sigma^{*}_{M}(\eta;\Lambda,N)\asymp\sqrt{M}\,e^{-\eta(\Lambda-\ell/2)}$
    with explicit multiplicative constants given by (i)–(ii).
2. Transfer matrices and Weyl function
For $z\in\mathbb{C}$, Equation˜1 can be written as
$Y^{\prime}(s)=-z\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,H(s)\,Y(s),$
since $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{-1}=-\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. For $s,t\in[0,\Lambda]$ define the transfer matrix $\Phi_{H}(s,t;z)$ by
$\partial_{s}\Phi_{H}(s,t;z)=-z\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,H(s)\,\Phi_{H}(s,t;z),\qquad\Phi_{H}(t,t;z)=I.$ (5)
Lemma 2.1 (Determinant one).
For each $z\in\mathbb{C}$ and $0\leq t\leq s\leq\Lambda$, $\det\Phi_{H}(s,t;z)=1$.
Proof.
Liouville’s formula for linear systems $Y^{\prime}=A(s)Y$ gives $\frac{d}{ds}\det\Phi=\operatorname{tr}(A(s))\det\Phi$. Here $A(s)=-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(s)$, so
$\frac{d}{ds}\det\Phi_{H}(s,t;z)=\operatorname{tr}\!\bigl(-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(s)\bigr)\det\Phi_{H}(s,t;z).$
Since $H(s)$ is symmetric and $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$ is skew-symmetric, $\operatorname{tr}(\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(s))=0$. Hence $\det\Phi_{H}(s,t;z)$ is constant in $s$, and equals $\det I=1$ at $s=t$; see [5]. ∎
Write the endpoint matrix
$\Phi_{H}(\Lambda;z):=\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}A_{H}(z)&B_{H}(z)\\ C_{H}(z)&D_{H}(z)\end{pmatrix}$.
Lemma 2.2 (Propagator bound).
Assume $H(s)\succeq 0$ and $\operatorname{tr}H(s)=1$ a.e. on $0,\Lambda]$. Then for all $z\in\mathbb{C}$ and $0\leq t\leq s\leq\Lambda$,
$\left\|\Phi_{H}(s,t;z)\right\|_{\mathrm{op}}\leq\exp\!\bigl(|z|(s-t)\bigr).$
Proof.
Fix $u\in\mathbb{C}^{2}$ with $\left\|u\right\|_{2}=1$ and set $Y(s)=\Phi_{H}(s,t;z)u$. Then $Y^{\prime}=-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(s)Y$. Moreover, $\left\|\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\right\|_{\mathrm{op}}=1$ and for PSD $2\times 2$ matrices one has $\left\|H(s)\right\|_{\mathrm{op}}\leq\operatorname{tr}H(s)=1$.
$\left\|Y^{\prime}(s)\right\|_{2}\leq|z|\,\left\|Y(s)\right\|_{2}\qquad\text{for a.e.\ }s.$
The function $s\mapsto\left\|Y(s)\right\|_{2}$ is absolutely continuous and satisfies $(\left\|Y\right\|_{2})^{\prime}\leq\left\|Y^{\prime}\right\|_{2}$ a.e.; Grönwall’s inequality then yields $\left\|Y(s)\right\|_{2}\leq e^{|z|(s-t)}$. Taking the supremum over $u$ yields the claim. ∎
2.1. Tail extension and Weyl coefficient
Extend $H$ to $0,\infty)$ by Equation˜2. For $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ there is (up to scalar) a unique solution of Equation˜1 that is square-integrable on $\Lambda,\infty)$ with respect to the $H$-weight, and the corresponding Weyl coefficient $m_{H,\Lambda}(z)$ is Herglotz–Nevanlinna; see, e.g., [9].
Theorem 2.3 (Transfer-matrix Weyl formula).
For any tail-extended Hamiltonian and $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$,
$m_{H,\Lambda}(z)=\frac{iA_{H}(z)-C_{H}(z)}{D_{H}(z)-iB_{H}(z)}.$ (6)
Moreover,
$D_{H}(z)-iB_{H}(z)\neq 0,\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}.$ (7)
Proof.
Write $\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}A&B\ C&D\end{pmatrix}$. On the tail $\Lambda,\infty)$ the system is $Y^{\prime}=-\frac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}Y$. For $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ the HH–square-integrable solution is proportional to $e^{iz(s-\Lambda)/2}u_{+}$, where $u_{+}=(1,i)^{\top}$. Thus the Weyl solution satisfies $Y(\Lambda)\parallel u_{+}$ (this is exactly the square-integrable tail mode for $\operatorname{Im}z>0$).
Write $Y(0)=(1,m)^{\top}$ and $Y(\Lambda)=\Phi_{H}(\Lambda,0;z)Y(0)$, hence
$u_{+}\ \parallel\ \binom{A+Bm}{C+Dm}$.
If $A+Bm=0$ then also $C+Dm=0$, so $(1,m)^{\top}\in\ker(\Phi_{H}(\Lambda,0;z))$, contradicting $\det\Phi_{H}(\Lambda,0;z)=1$ by Lemma˜2.1. Thus $A+Bm\neq 0$, and taking ratios gives
$i=\frac{C+Dm}{A+Bm}\quad\Rightarrow\quad i(A+Bm)=C+Dm,$
which rearranges to Equation˜6.
For Equation˜7, suppose DiB=0D-iB=0 at some $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. Then $D=iB$. From $i(A+Bm)=C+Dm$ we cancel the $m$-terms and obtain $iA=C$. Then
$\det\Phi_{H}(\Lambda,0;z)=AD-BC=A(iB)-B(iA)=0,$
contradicting $\det\Phi_{H}(\Lambda,0;z)=1$. Hence $D-iB\neq 0$ on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. ∎
2.2. A perturbation formula for the Weyl function as a spectral pairing
Lemma 2.4 (J–symplecticity of transfer matrices).
For every $z\in\mathbb{C}$ and $0\leq t\leq s\leq\Lambda$ one has
$\Phi_{H}(s,t;z)^{\top}\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Phi_{H}(s,t;z)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$.
Proof.
Set $W(s):=\Phi_{H}(s,t;z)^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Phi_{H}(s,t;z)$. Differentiating and using
$\partial_{s}\Phi_{H}(s,t;z)=-z\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(s)\Phi_{H}(s,t;z),$
together with $H(s)^{\top}=H(s)$ and $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{\top}=-\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$, we obtain
$\partial_{s}W(s)$ $=(\partial_{s}\Phi_{H})^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Phi_{H}+\Phi_{H}^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}(\partial_{s}\Phi_{H})$
$=z\,\Phi_{H}^{\top}H(s)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{2}\Phi_{H}-z\,\Phi_{H}^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{2}H(s)\,\Phi_{H}\;=\;0,$
since $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{2}=-I$. Hence $W(s)$ is constant and $W(s)\equiv W(t)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. ∎
Theorem 2.5 (First variation of m as a rank-one quadratic form).
Fix $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and let $m(z)=m_{H,\Lambda}(z)$. Let $Y(\cdot;z)$ be the Weyl solution on $0,\infty)$ normalized by $Y(0;z)=\binom{1}{m(z)}$. Let $\Delta H\in L^{\infty}(0,\Lambda;\mathbb{R}^{2\times 2})$ be traceless a.e., and assume $H_{\varepsilon}:=H+\varepsilon\Delta H$ remains nonnegative for $|\varepsilon|$ small. Then $\varepsilon\mapsto m_{H_{\varepsilon},\Lambda}(z)$ is differentiable at $\varepsilon=0$ and
$Dm_{H,\Lambda}(z)[\Delta H]=z\int_{0}^{\Lambda}Y(s;z)^{\top}\,\Delta H(s)\,Y(s;z)\,ds.$ (8)
Proof.
Let $u_{+}:=\binom{1}{i}$ and write $\Phi(\cdot,\cdot):=\Phi_{H}(\cdot,\cdot;z)$. Set $\Phi:=\Phi(\Lambda,0)$ and $g:=\Phi^{-1}u_{+}\in\mathbb{C}^{2}$. Since $m(z)$ is finite on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, we have $g_{1}\neq 0$ and $m(z)=g_{2}/g_{1}$ by Theorem 2.3. Define $y(s):=\Phi(s,0)g$, so $y(\Lambda)=u_{+}$, and note that
$Y(s;z)=\frac{1}{g_{1}}\,y(s),\qquad 0\leq s\leq\Lambda.$
For $|\varepsilon|$ small let $\Phi_{\varepsilon}(\cdot,\cdot)$ be the transfer matrix for $H_{\varepsilon}$ at the same $z$ and set $\Phi_{\varepsilon}:=\Phi_{\varepsilon}(\Lambda,0)$ and $g_{\varepsilon}:=\Phi_{\varepsilon}^{-1}u_{+}$. Then $m_{H_{\varepsilon},\Lambda}(z)=(g_{\varepsilon})_{2}/(g_{\varepsilon})_{1}$.
Differentiate the identity $\Phi_{\varepsilon}g_{\varepsilon}=u_{+}$ at $\varepsilon=0$ to obtain $\dot{g}=-\Phi^{-1}\dot{\Phi}\,g$ (dot = $\partial_{\varepsilon}|_{\varepsilon=0}$). The standard variation-of-constants formula gives; see [5].; see [5].
$\dot{\Phi}=-z\int_{0}^{\Lambda}\Phi(\Lambda,s)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,\Phi(s,0)\,ds.$ (9)
Indeed, $\partial_{s}\Phi_{\varepsilon}(s,0)=-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H_{\varepsilon}(s)\Phi_{\varepsilon}(s,0)$; differentiating in $\varepsilon$ and solving the resulting inhomogeneous linear system for $\dot{\Phi}(\cdot,0)$ yields (9).
Using $\dot{g}=-\Phi^{-1}\dot{\Phi}\,g$ and $\Phi^{-1}\Phi(\Lambda,s)=\Phi(s,0)^{-1}$, (9) yields
$\dot{g}=z\int_{0}^{\Lambda}\Phi(s,0)^{-1}\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,y(s)\,ds.$
Now $m=g_{2}/g_{1}$, so
$\dot{m}=\frac{g_{1}\dot{g}_{2}-g_{2}\dot{g}_{1}}{g_{1}^{2}}.$
Introduce the bilinear form $\omega(a,b):=a_{1}b_{2}-a_{2}b_{1}$. Since $\omega(a,b)=-a^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}b$, we have $g_{1}\dot{g}_{2}-g_{2}\dot{g}_{1}=\omega(g,\dot{g})=-g^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\dot{g}$ and therefore
$\dot{m}=-\frac{1}{g_{1}^{2}}\,g^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\dot{g}=-\frac{z}{g_{1}^{2}}\int_{0}^{\Lambda}g^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Phi(s,0)^{-1}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,y(s)\,ds.$
Since $\Phi(s,0)^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Phi(s,0)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$ by Lemma 2.4, we have
$\Phi(s,0)^{-\top}\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Phi(s,0)^{-1}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$,
hence $g^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Phi(s,0)^{-1}=y(s)^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. Thus
$\dot{m}=-\frac{z}{g_{1}^{2}}\int_{0}^{\Lambda}y(s)^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,y(s)\,ds=\frac{z}{g_{1}^{2}}\int_{0}^{\Lambda}y(s)^{\top}\,\Delta H(s)\,y(s)\,ds.$
Recalling $Y=y/g_{1}$ on $0,\Lambda]$ gives (8). ∎
Corollary 2.6 (Two-kernel form along traceless directions).
Let $q\in L^{\infty}(0,\Lambda;\mathbb{C})$ and $\Delta H=\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}$. Write the normalized Weyl solution in the $u_{\pm}$ basis as
$Y(s;z)=a(s;z)\,u_{+}+b(s;z)\,u_{-},\qquad u_{\pm}=\binom{1}{\pm i}$.
Then for every $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$,
$Dm_{H,\Lambda}(z)[q]=2z\int_{0}^{\Lambda}\big(q(s)\,a(s;z)^{2}+\overline{q(s)}\,b(s;z)^{2}\big)\,ds.$
In particular, at the free Hamiltonian $H_{0}=\tfrac{1}{2}I$ one has $b(\cdot;z)\equiv 0$ and $Dm_{H_{0},\Lambda}(z)[q]=2z\int_{0}^{\Lambda}q(s)e^{izs}\,ds$.
Proof.
The identities $\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,u_{+}=q\,u_{-}$ and $\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,u_{-}=\overline{q}\,u_{+}$ are immediate matrix multiplications. Since $u_{+}^{\top}u_{-}=2$ and $u_{\pm}^{\top}u_{\pm}=0$, we obtain
$Y^{\top}\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,Y=(au_{+}+bu_{-})^{\top}\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,(au_{+}+bu_{-})=2q\,a^{2}+2\overline{q}\,b^{2}.$
Insert this into (8). The free case follows from the explicit Weyl solution $Y(s;z)=e^{izs/2}u_{+}$ for $H_{0}$. ∎
Remark 2.7 (From pairing to operator language).
For each fixed $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, (8) shows that $\Delta H\mapsto Dm_{H,\Lambda}(z)[\Delta H]$ is the bounded linear functional given by pairing $\Delta H$ with the rank-one field $s\mapsto Y(s;z)Y(s;z)^{\top}$. When $z=x+i\eta$ with $\eta>0$, the decay of the Weyl solution on the free tail forces $Y(s;z)$ to carry an exponential factor $\sim e^{-\eta(s-\Lambda)/2}$ for $s\geq\Lambda$, and the depth attenuation in the Jacobian is visible directly from this quadratic form.
3. Base linearization at the free Hamiltonian
Let $H_{0}\equiv\tfrac{1}{2}I$ on $0,\Lambda]$ (and tail). Define the eigenvectors
$u_{+}=(1,i)^{\top},\qquad u_{-}=(1,-i)^{\top}.$
We will often expand solutions in the $\{u_{+},u_{-}\}$ basis.
Lemma 3.1 (J-eigenvectors and X-matrix action).
One has $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}u_{+}=-iu_{+}$ and $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}u_{-}=iu_{-}$. Moreover, for all $q\in\mathbb{C}$,
$\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,u_{+}=q\,u_{-},\qquad\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}\,u_{-}=\overline{q}\,u_{+}.$
Proof.
Compute $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}(1,i)^{\top}=(-i,1)^{\top}=-i(1,i)^{\top}$. Similarly, $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}(1,-i)^{\top}=(i,1)^{\top}=i(1,-i)^{\top}$. For the $\begin{pmatrix}\operatorname{Re}\cdot&\operatorname{Im}\cdot\\ \operatorname{Im}\cdot&-\operatorname{Re}\cdot\end{pmatrix}$ identities, write $q=a+ib$; then $\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}=\begin{pmatrix}a&b\\ b&-a\end{pmatrix}$ and multiply by $u_{\pm}$:
$\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}u_{+}$ $=\binom{a+bi}{\,b-ai}$ (10)
$=(a+ib)\binom{1}{-i}$ (11)
$=q\,u_{-},$ (12)
$\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}u_{-}$ $=\binom{a-bi}{\,b+ai}$ (13)
$=\overline{q}\,\binom{1}{i}$ (14)
$=\overline{q}\,u_{+}.$ (15)
Lemma 3.2 (Free propagator and base Weyl value).
For $H_{0}\equiv\tfrac{1}{2}I$ and any $z\in\mathbb{C}$,
$\Phi_{H_{0}}(s,t;z)=\exp\!\left(-\frac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}(s-t)\right)=\begin{pmatrix}\cos(\tfrac{z}{2}(s-t))&\sin(\tfrac{z}{2}(s-t))\\ -\sin(\tfrac{z}{2}(s-t))&\cos(\tfrac{z}{2}(s-t))\end{pmatrix}$.
For $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, one has $m_{H_{0},\Lambda}(z)\equiv i$ and hence $v_{H_{0},\Lambda}(z)\equiv 0$. Moreover,
$D_{H_{0}}(z)-iB_{H_{0}}(z)=e^{-iz\Lambda/2},\qquad\left|D_{H_{0}}(z)-iB_{H_{0}}(z)\right|=e^{(\operatorname{Im}z)\Lambda/2}.$ (16)
Proof.
With $H_{0}=\tfrac{1}{2}I$, Equation˜5 is $\Phi^{\prime}=-\frac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Phi$ and $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{2}=-I$; thus $\Phi=\exp(-(z/2)\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}(s-t))=\cos(\tfrac{z}{2}(s-t))I-\sin(\tfrac{z}{2}(s-t))\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. At $s=\Lambda,t=0$, $A_{0}=D_{0}=\cos(z\Lambda/2)$, $B_{0}=\sin(z\Lambda/2)$, $C_{0}=-\sin(z\Lambda/2)$; hence
$iA_{0}-C_{0}=i\cos(z\Lambda/2)+\sin(z\Lambda/2)=i(\cos(z\Lambda/2)-i\sin(z\Lambda/2))=i(D_{0}-iB_{0}).$
By Theorem˜2.3, $m\equiv i$ and hence $v\equiv 0$. Finally $D_{0}-iB_{0}=e^{-iz\Lambda/2}$ and $\left|e^{-iz\Lambda/2}\right|=e^{(\operatorname{Im}z)\Lambda/2}$. ∎
3.1. First variation at $H_{0}$
Lemma 3.3 (Duhamel formula at the base).
Let $\Delta H\in L^{\infty}(0,\Lambda;\mathbb{R}^{2\times 2}_{\mathrm{sym}})$. Then for each $z\in\mathbb{C}$,
$D\Phi_{H_{0}}(\Lambda,0;z)[\Delta H]=-z\int_{0}^{\Lambda}\Phi_{H_{0}}(\Lambda,s;z)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,\Phi_{H_{0}}(s,0;z)\,ds.$
Proof.
Set $\Phi_{\varepsilon}$ to be the transfer matrix for $H_{\varepsilon}:=H_{0}+\varepsilon\Delta H$. Differentiate the ODE $\partial_{s}\Phi_{\varepsilon}=-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H_{\varepsilon}(s)\Phi_{\varepsilon}$ in $\varepsilon$ and evaluate at $\varepsilon=0$:
$\partial_{s}(\partial_{\varepsilon}\Phi_{\varepsilon})|_{\varepsilon=0}=-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H_{0}(s)\,(\partial_{\varepsilon}\Phi_{\varepsilon})|_{\varepsilon=0}-z\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\Delta H(s)\Phi_{H_{0}}(s,0;z),$
with $(\partial_{\varepsilon}\Phi_{\varepsilon})(0,0)=0$. Variation-of-constants yields the stated formula; see [5]. ∎
3.2. A base differentiation lemma
Lemma 3.4 (Differentiating m and v at the free point).
Fix $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and write $\Phi_{0}:=\Phi_{H_{0}}(\Lambda,0;z)=\begin{pmatrix}A_{0}&B_{0}\\ C_{0}&D_{0}\end{pmatrix}$. Let $\delta\Phi=\begin{pmatrix}\delta A&\delta B\\ \delta C&\delta D\end{pmatrix}$ be a first variation of $\Phi(\Lambda,0;z)$. Then the induced first variations $\delta m$ and $\delta v$ at the free point satisfy
$\delta m=\frac{i(\delta A+i\delta B)-(\delta C+i\delta D)}{D_{0}-iB_{0}},$ (17)
and, since $v=\frac{m-i}{m+i}$,
$\delta v=\frac{\delta m}{2i}=\frac{i(\delta A+i\delta B)-(\delta C+i\delta D)}{2i\,(D_{0}-iB_{0})}.$ (18)
In particular, writing $u_{+}=(1,i)^{\top}$ and $u_{-}=(1,-i)^{\top}$, if $\delta\Phi\,u_{+}$ is proportional to $u_{-}$, i.e.
$\delta\Phi\,u_{+}=\beta\,u_{-}\qquad\text{for some }\beta\in\mathbb{C},$ (19)
then necessarily $\delta A+i\delta B=\beta$ and $\delta C+i\delta D=-i\beta$, and (18) simplifies to
$\delta v=\frac{\beta}{D_{0}-iB_{0}}.$ (20)
Proof.
At general $H$, $m=\frac{iA-C}{D-iB}$. Differentiating and evaluating at the free point where $m_{0}=i$ yields
$\delta m=\frac{(i\delta A-\delta C)(D_{0}-iB_{0})-(iA_{0}-C_{0})(\delta D-i\delta B)}{(D_{0}-iB_{0})^{2}}.$
By Lemma˜3.2, $iA_{0}-C_{0}=i(D_{0}-iB_{0})$, so cancelling a common factor $D_{0}-iB_{0}$ gives
$\delta m=\frac{i\delta A-\delta C-i(\delta D-i\delta B)}{D_{0}-iB_{0}}=\frac{i(\delta A+i\delta B)-(\delta C+i\delta D)}{D_{0}-iB_{0}},$
which is (17).
Since $v(m)=\frac{m-i}{m+i}$, we have $v^{\prime}(m)=\frac{2i}{(m+i)^{2}}$, hence $v^{\prime}(i)=\frac{1}{2i}$ and therefore $\delta v=\frac{\delta m}{2i}$, giving (18).
Finally, if (19) holds, then
$\binom{\delta A+i\delta B}{\delta C+i\delta D}=\delta\Phi\,u_{+}=\beta\binom{1}{-i}$,
so $\delta A+i\delta B=\beta$ and $\delta C+i\delta D=-i\beta$, and (20) follows by substitution into (18). ∎
3.3. Continuum first variation: Laplace/Fourier form
Theorem 3.5 (Continuum base linearization and depth weight).
Let $H_{0}\equiv\tfrac{1}{2}I$ and let $q\in L^{1}(0,\Lambda;\mathbb{C})$. Define the linear functional
$\mathcal{L}_{z}(q):=-iz\int_{0}^{\Lambda}q(s)\,e^{izs}\,ds,\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}.$
For $q\in L^{1}(0,\Lambda;\mathbb{C})\cap L^{\infty}(0,\Lambda;\mathbb{C})$, the Fréchet derivative of $v_{H_{0},\Lambda}$ in the direction $\Delta H_{q}(s):=\begin{pmatrix}\operatorname{Re}q(s)&\operatorname{Im}q(s)\\ \operatorname{Im}q(s)&-\operatorname{Re}q(s)\end{pmatrix}$ is given by
$Dv_{H_{0},\Lambda}(z)[q]=\mathcal{L}_{z}(q).$ (21)
Moreover, $\mathcal{L}_{z}$ extends uniquely (by continuity) to a bounded linear functional on all of $L^{1}(0,\Lambda;\mathbb{C})$, and we use the same notation $Dv_{H_{0},\Lambda}(z)[q]:=\mathcal{L}_{z}(q)$ for this $L^{1}$-continuous extension.
In particular,
$\left|Dv_{H_{0},\Lambda}(x+i\eta)[q]\right|\leq|z|\int_{0}^{\Lambda}|q(s)|\,e^{-\eta s}\,ds.$
If $q$ is supported in $s_{\star},\Lambda]$ then
$\left|Dv_{H_{0},\Lambda}(x+i\eta)[q]\right|\leq|z|\,e^{-\eta s_{\star}}\,\left\|q\right\|_{L^{1}(0,\Lambda)}.$
Proof.
We first establish (21) for $q\in L^{\infty}(0,\Lambda)$ and then extend by density to $q\in L^{1}$. Let $R(t):=\Phi_{H_{0}}(t,0;z)=\exp(-\tfrac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}t)$. By Lemma˜3.3 (which applies for $\Delta H\in L^{\infty}$),
$\delta\Phi(\Lambda,0;z)=-z\int_{0}^{\Lambda}R(\Lambda-s)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\begin{pmatrix}\operatorname{Re}q(s)&\operatorname{Im}q(s)\\ \operatorname{Im}q(s)&-\operatorname{Re}q(s)\end{pmatrix}\,R(s)\,ds.$
Using Lemma˜3.1, $R(s)u_{+}=e^{izs/2}u_{+}$ and also $R(\Lambda-s)u_{-}=e^{-iz(\Lambda-s)/2}u_{-}$, hence
$R(\Lambda-s)\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}\operatorname{Re}q(s)&\operatorname{Im}q(s)\\ \operatorname{Im}q(s)&-\operatorname{Re}q(s)\end{pmatrix}R(s)u_{+}$ $=i\,q(s)\,e^{-iz(\Lambda-s)/2}e^{izs/2}\,u_{-}$
$=i\,q(s)\,e^{-iz\Lambda/2}e^{izs}\,u_{-}.$
Therefore
$\delta\Phi(\Lambda,0;z)u_{+}=-iz\,e^{-iz\Lambda/2}\left(\int_{0}^{\Lambda}q(s)e^{izs}\,ds\right)u_{-}.$
So $\delta\Phi\,u_{+}=\beta u_{-}$ with $\beta=-iz\,e^{-iz\Lambda/2}\int_{0}^{\Lambda}q(s)e^{izs}\,ds$. Apply Lemma˜3.4 and Equation˜16: since $D_{0}-iB_{0}=e^{-iz\Lambda/2}$, we obtain
$\delta v=\frac{\beta}{D_{0}-iB_{0}}=-iz\int_{0}^{\Lambda}q(s)e^{izs}\,ds,$
proving (21) for $q\in L^{\infty}$.
For general $q\in L^{1}$, choose $q_{n}\in L^{\infty}\cap L^{1}$ with $q_{n}\to q$ in $L^{1}$. The right-hand side of (21) is continuous in $L^{1}$ by $\bigl|\int_{0}^{\Lambda}(q_{n}-q)(s)e^{izs}\,ds\bigr|\leq\|q_{n}-q\|_{L^{1}}$. Thus the functionals $\mathcal{L}_{z}(q_{n})=-iz\int_{0}^{\Lambda}q_{n}(s)e^{izs}\,ds$ converge to $\mathcal{L}_{z}(q)$, and $\mathcal{L}_{z}$ defines a unique bounded linear extension from $L^{1}\cap L^{\infty}$ to all of $L^{1}$. We adopt the notation $Dv_{H_{0},\Lambda}(z)[q]:=\mathcal{L}_{z}(q)$ for this $L^{1}$-continuous extension. The bounds follow from $\left|e^{izs}\right|=e^{-\eta s}$. ∎
3.4. Quadratic remainder bounds in $L^{1}$
This subsection provides quantitative remainder estimates that complement Theorem˜3.5. The key input is a correct Volterra integral equation for $\Phi_{H}(\cdot,0;z)$, which holds for $L^{1}$ coefficients.
Lemma 3.6 (Volterra integral equation).
Let $H\in L^{1}(0,\Lambda;\mathbb{R}^{2\times 2}_{\mathrm{sym}})$ and fix $z\in\mathbb{C}$. Then $X(s):=\Phi_{H}(s,0;z)$ satisfies
$X(s)=I-z\int_{0}^{s}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(u)\,X(u)\,du,\qquad 0\leq s\leq\Lambda.$ (22)
Equivalently, for any $0\leq t\leq s\leq\Lambda$,
$\Phi_{H}(s,t;z)=I-z\int_{t}^{s}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}H(u)\,\Phi_{H}(u,t;z)\,du.$ (23)
Proof.
Integrate the differential equation in Equation˜5 from 0 to $s$ and use $\Phi_{H}(0,0;z)=I$. The statement for $\Phi_{H}(s,t;z)$ follows by the cocycle identity $\Phi_{H}(u,0)=\Phi_{H}(u,t)\Phi_{H}(t,0)$ and the same integration on $[t,s]$; see [5]. ∎
Lemma 3.7 (Dyson expansion and a quadratic remainder for Phi).
Fix $z\in\mathbb{C}$ and write $\Phi_{\Delta}:=\Phi_{H_{0}+\Delta H}(\Lambda,0;z)$, $\Phi_{0}:=\Phi_{H_{0}}(\Lambda,0;z)$. Assume $\Delta H\in L^{1}(0,\Lambda;\mathbb{R}^{2\times 2}_{\mathrm{sym}})$ and $\|\Delta H\|_{L^{\infty}(0,\Lambda;\mathrm{op})}:=\operatorname*{ess\,sup}_{s\in(0,\Lambda)}\|\Delta H(s)\|_{\mathrm{op}}\leq\tfrac{1}{4}$. Then $H_{0}+\Delta H\succeq\tfrac{1}{4}I$ a.e. and $\Phi_{\Delta}$ exists. Moreover, defining the first-order term
$L(\Delta H):=-z\int_{0}^{\Lambda}\Phi_{0}(\Lambda,s;z)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,\Phi_{0}(s,0;z)\,ds,$
one has the second-order estimate
$\|\Phi_{\Delta}-\Phi_{0}-L(\Delta H)\|_{\mathrm{op}}\ \leq\ |z|^{2}\,\exp\!\bigl(4|z|\Lambda\bigr)\,\|\Delta H\|_{L^{1}(0,\Lambda)}^{2}.$ (24)
Proof.
Step 1: uniform propagator bounds. Since $\|H_{0}\|_{\mathrm{op}}=1/2$ and $\|H_{0}+\Delta H\|_{\mathrm{op}}\leq 1/2+1/4=3/4$ a.e., the Grönwall argument in Lemma˜2.2 (using $\|H\|_{\mathrm{op}}$ in place of $\operatorname{tr}H$) yields, for $0\leq t\leq s\leq\Lambda$,
$\|\Phi_{H_{0}+\Delta H}(s,t;z)\|_{\mathrm{op}}\leq\exp(|z|(s-t))\leq e^{|z|\Lambda},\qquad\|\Phi_{H_{0}}(s,t;z)\|_{\mathrm{op}}\leq e^{|z|\Lambda}.$ (25)
Step 2: an exact Duhamel identity. Let $\Phi_{\Delta}(s,t):=\Phi_{H_{0}+\Delta H}(s,t;z)$ and $\Phi_{0}(s,t):=\Phi_{H_{0}}(s,t;z)$. A standard variation-of-constants argument (equivalently, subtract the integral equations (23) for $H_{0}+\Delta H$ and $H_{0}$ and see [5]) gives, for $0\leq t\leq s\leq\Lambda$, solve for $\Phi_{\Delta}-\Phi_{0}$) gives, for $0\leq t\leq s\leq\Lambda$,
$\Phi_{\Delta}(s,t)-\Phi_{0}(s,t)=-z\int_{t}^{s}\Phi_{\Delta}(s,u)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(u)\,\Phi_{0}(u,t)\,du.$ (26)
In particular, using (25),
$\sup_{0\leq t\leq s\leq\Lambda}\|\Phi_{\Delta}(s,t)-\Phi_{0}(s,t)\|_{\mathrm{op}}\leq|z|\,e^{2|z|\Lambda}\,\|\Delta H\|_{L^{1}(0,\Lambda)}.$ (27)
Step 3: second-order remainder at $(\Lambda,0)$. Write (26) at $(\Lambda,0)$, and insert $\Phi_{\Delta}(\Lambda,u)=\Phi_{0}(\Lambda,u)+(\Phi_{\Delta}(\Lambda,u)-\Phi_{0}(\Lambda,u))$:
$\Phi_{\Delta}(\Lambda,0)-\Phi_{0}(\Lambda,0)$ $=-z\int_{0}^{\Lambda}\Phi_{0}(\Lambda,u)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(u)\,\Phi_{0}(u,0)\,du$
$\quad-z\int_{0}^{\Lambda}(\Phi_{\Delta}(\Lambda,u)-\Phi_{0}(\Lambda,u))\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(u)\,\Phi_{0}(u,0)\,du.$
The first term is exactly $L(\Delta H)$. Bound the second term using (25) and (27):
$\|\Phi_{\Delta}-\Phi_{0}-L(\Delta H)\|_{\mathrm{op}}$ $\leq|z|\int_{0}^{\Lambda}\|\Phi_{\Delta}(\Lambda,u)-\Phi_{0}(\Lambda,u)\|_{\mathrm{op}}\,\|\Delta H(u)\|_{\mathrm{op}}\,\|\Phi_{0}(u,0)\|_{\mathrm{op}}\,du$
$\leq|z|\,(|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}})\,e^{|z|\Lambda}\int_{0}^{\Lambda}\|\Delta H(u)\|_{\mathrm{op}}\,du$
$\leq|z|^{2}\,e^{3|z|\Lambda}\,\|\Delta H\|_{L^{1}}^{2}.$
Inflate $e^{3|z|\Lambda}$ to $e^{4|z|\Lambda}$ to obtain (24). ∎
Theorem 3.8 (Quadratic remainder for m and v at the free point).
Fix $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. Let $\Delta H\in L^{1}(0,\Lambda;\mathbb{R}^{2\times 2}_{\mathrm{sym}})$ be traceless a.e. and satisfy $\|\Delta H\|_{L^{\infty}(0,\Lambda;\mathrm{op})}:=\operatorname*{ess\,sup}_{s\in(0,\Lambda)}\|\Delta H(s)\|_{\mathrm{op}}\leq\tfrac{1}{4}$. Define $m_{\Delta}(z):=m_{H_{0}+\Delta H,\Lambda}(z)$ and $v_{\Delta}(z):=v_{H_{0}+\Delta H,\Lambda}(z)$.
Set
$r_{m}(z,\Lambda):=\frac{1}{8|z|}\,e^{-2|z|\Lambda}\,e^{(\operatorname{Im}z)\Lambda/2}.$ (28)
Assume in addition that
$\|\Delta H\|_{L^{1}(0,\Lambda)}\leq r_{m}(z,\Lambda).$ (29)
Then there exists an explicit constant $C=C(z,\Lambda)$ such that
$\left|m_{\Delta}(z)-i-Dm_{H_{0},\Lambda}(z)[\Delta H]\right|\ \leq\ C\,\|\Delta H\|_{L^{1}(0,\Lambda)}^{2},$ (30)
and
$\left|v_{\Delta}(z)-Dv_{H_{0},\Lambda}(z)[\Delta H]\right|\ \leq\ C\,\|\Delta H\|_{L^{1}(0,\Lambda)}^{2}.$ (31)
One may take
$C(z,\Lambda):=64\,|z|^{2}\,\exp\!\bigl(6|z|\Lambda\bigr)\,e^{-(\operatorname{Im}z)\Lambda/2}.$
Proof.
Let $\Phi_{\Delta}=\begin{pmatrix}A_{\Delta}&B_{\Delta}\\ C_{\Delta}&D_{\Delta}\end{pmatrix}$ and $\Phi_{0}=\begin{pmatrix}A_{0}&B_{0}\\ C_{0}&D_{0}\end{pmatrix}$. Define $n_{\Delta}:=iA_{\Delta}-C_{\Delta}$ and $d_{\Delta}:=D_{\Delta}-iB_{\Delta}$ so that $m_{\Delta}=n_{\Delta}/d_{\Delta}$ by Theorem˜2.3. At the base, $n_{0}=id_{0}$ with $d_{0}=e^{-iz\Lambda/2}$ by Lemma˜3.2.
By Lemma˜3.7 with $L(\Delta H)=D\Phi_{H_{0}}[\Delta H]$,
$\|\Phi_{\Delta}-\Phi_{0}-L(\Delta H)\|_{\mathrm{op}}\leq|z|^{2}e^{4|z|\Lambda}\|\Delta H\|_{L^{1}}^{2}.$
Hence each entry of $\Phi_{\Delta}-\Phi_{0}-L(\Delta H)$ is bounded by the same right-hand side, and therefore
$\left|\delta n-(i\,\delta A^{(1)}-\delta C^{(1)})\right|\ +\ \left|\delta d-(\delta D^{(1)}-i\,\delta B^{(1)})\right|\ \leq\ 4|z|^{2}e^{4|z|\Lambda}\|\Delta H\|_{L^{1}}^{2},$ (32)
where $\delta n:=n_{\Delta}-n_{0}$, $\delta d:=d_{\Delta}-d_{0}$, and $\delta A^{(1)},\dots,\delta D^{(1)}$ denote the entries of $L(\Delta H)$.
Next we ensure a uniform denominator margin. From (26) at $(\Lambda,0)$ and (25) (with $e^{|z|\Lambda}$ for both $\Phi_{\Delta}$ and $\Phi_{0}$),
$\|\Phi_{\Delta}(\Lambda,0)-\Phi_{0}(\Lambda,0)\|_{\mathrm{op}}\leq|z|\,e^{2|z|\Lambda}\,\|\Delta H\|_{L^{1}}.$
Since entrywise changes are bounded by operator-norm changes in $2\times 2$, we obtain
$|d_{\Delta}-d_{0}|\leq|D_{\Delta}-D_{0}|+|B_{\Delta}-B_{0}|\leq 2\|\Phi_{\Delta}-\Phi_{0}\|_{\mathrm{op}}\leq 2|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}.$
Under (29) this implies $|d_{\Delta}-d_{0}|\leq\tfrac{1}{4}|d_{0}|$ because $|d_{0}|=e^{(\operatorname{Im}z)\Lambda/2}$ by Equation˜16. Hence
$|d_{\Delta}|\geq\tfrac{3}{4}\,|d_{0}|=\tfrac{3}{4}e^{(\operatorname{Im}z)\Lambda/2}.$ (33)
Now write the exact identity
$m_{\Delta}-i=\frac{n_{\Delta}-id_{\Delta}}{d_{\Delta}}=\frac{(n_{0}+\delta n)-i(d_{0}+\delta d)}{d_{\Delta}}=\frac{\delta n-i\delta d}{d_{0}}+\frac{\delta n-i\delta d}{d_{\Delta}}\Bigl(\frac{d_{0}-d_{\Delta}}{d_{0}}\Bigr).$
Let
$n^{(1)}:=i\,\delta A^{(1)}-\delta C^{(1)},\qquad d^{(1)}:=\delta D^{(1)}-i\,\delta B^{(1)}.$
Since $n_{0}-id_{0}=0$, the Fréchet derivative of $m$ at $H_{0}$ is
$Dm_{H_{0},\Lambda}(z)[\Delta H]=\frac{n^{(1)}-i\,d^{(1)}}{d_{0}},$ (34)
by Lemma˜3.4 (equivalently, by differentiating $m=n/d$ at $(n_{0},d_{0})$).
Subtracting (34) from the exact identity yields
$m_{\Delta}-i-Dm_{H_{0},\Lambda}(z)[\Delta H]=\frac{(\delta n-i\delta d)-(n^{(1)}-i\,d^{(1)})}{d_{0}}+\frac{\delta n-i\delta d}{d_{\Delta}}\Bigl(\frac{d_{0}-d_{\Delta}}{d_{0}}\Bigr).$
The first term is quadratic by (32). For the second term, note that $\left|\delta n-i\delta d\right|\leq 4|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ by the Lipschitz bound $\|\Phi_{\Delta}-\Phi_{0}\|_{\mathrm{op}}\leq|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ and the definitions of $n_{\Delta},d_{\Delta}$, while $\left|d_{0}-d_{\Delta}\right|\leq 2|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ and $\left|d_{\Delta}\right|^{-1}\leq\tfrac{4}{3}e^{-(\operatorname{Im}z)\Lambda/2}$ by (33). Combining these bounds with (32) and inflating constants yields (30) with the stated $C(z,\Lambda)$. Finally, the Cayley transform $w\mapsto(w-i)/(w+i)$ has uniformly bounded first and second derivatives on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (e.g. by direct differentiation and $|w+i|\geq 1$ for $w\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$), so the same quadratic bound transfers from $m$ to $v$, giving (31). ∎
Theorem 3.9 (Robust depth attenuation near H0 (two-kernel form)).
Fix $\Lambda>0$ and $z=x+i\eta\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. Assume the free tail $H\equiv\tfrac{1}{2}I$ on $\Lambda,\infty)$. Let $H$ be trace-normed on $0,\Lambda]$ and write it a.e. as
$H(s)=\tfrac{1}{2}I+\begin{pmatrix}\operatorname{Re}p(s)&\operatorname{Im}p(s)\\ \operatorname{Im}p(s)&-\operatorname{Re}p(s)\end{pmatrix},\qquad p\in L^{1}(0,\Lambda;\mathbb{C}),$
with positivity margin $\|p\|_{L^{\infty}(0,\Lambda)}\leq\tfrac{1}{2}-\varepsilon$ for some $\varepsilon\in(0,\tfrac{1}{2})$.
There exist $\delta>0$ and constants $C_{1},C_{2}$ (depending only on $z,\Lambda,\varepsilon$) such that if
$\|p\|_{L^{1}(0,\Lambda)}\leq\delta,$
then for every traceless perturbation $\Delta H(s)=\begin{pmatrix}\operatorname{Re}q(s)&\operatorname{Im}q(s)\\ \operatorname{Im}q(s)&-\operatorname{Re}q(s)\end{pmatrix}$ with $q\in L^{1}(0,\Lambda)\cap L^{\infty}(0,\Lambda)$ the Fréchet derivative (real-linear in $q$) satisfies
$Dv_{H,\Lambda}(z)[q]=-iz\int_{0}^{\Lambda}\Bigl(q(s)\,K^{(1)}_{H}(s;z)+\overline{q(s)}\,K^{(2)}_{H}(s;z)\Bigr)\,ds,$ (35)
where $K^{(1)}_{H},K^{(2)}_{H}\in L^{\infty}(0,\Lambda)$ obey
$|K^{(1)}_{H}(s;z)|\leq C_{1}\,e^{-\eta s}\qquad(0\leq s\leq\Lambda),\qquad\|K^{(2)}_{H}(\cdot;z)\|_{L^{\infty}(0,\Lambda)}\leq C_{2}\,\|p\|_{L^{1}(0,\Lambda)}^{2}.$
Moreover, for $H=H_{0}$ one has $K^{(1)}_{H_{0}}(s;z)=e^{izs}$ and $K^{(2)}_{H_{0}}\equiv 0$.
In particular, if $\operatorname{supp}q\subset[s_{\star},\Lambda]$, then
$|Dv_{H,\Lambda}(z)[q]|\leq|z|\Bigl(C_{1}e^{-\eta s_{\star}}+C_{2}\|p\|_{L^{1}}^{2}\Bigr)\,\|q\|_{L^{1}(0,\Lambda)}.$
Proof.
Step 0: $u_{\pm}$ basis and ratio formula for $v$. Let
$u_{+}=(1,i)^{\top},\quad u_{-}=(1,-i)^{\top},\quad P:=\begin{pmatrix}1&1\\ i&-i\end{pmatrix}=(u_{+}\ \ u_{-}),\quad\widetilde{\Phi}_{H}:=P^{-1}\Phi_{H}P.$
Any vector $y=au_{+}+bu_{-}$ satisfies $y_{1}=a+b$, $y_{2}=i(a-b)$, hence $m=y_{2}/y_{1}=i\frac{a-b}{a+b}$ and therefore
$v=\frac{m-i}{m+i}=-\frac{b}{a}.$
With the free tail on $\Lambda,\infty)$ and $\eta>0$, the unique HH–square-integrable mode at $\infty$ is proportional to $u_{+}$, so at $s=\Lambda$ the Weyl solution has $b(\Lambda)=0$. Writing $(a(\Lambda),b(\Lambda))^{\top}=\widetilde{\Phi}_{H}(\Lambda,0;z)\,(a(0),b(0))^{\top}$ and denoting $\widetilde{\Phi}_{H}(\Lambda,0;z)=:\begin{pmatrix}\tilde{a}&\tilde{b}\\ \tilde{c}&\tilde{d}\end{pmatrix}$, we have
$0=b(\Lambda)=\tilde{c}\,a(0)+\tilde{d}\,b(0)\quad\Rightarrow\quad v_{H,\Lambda}(z)=-\frac{b(0)}{a(0)}=\frac{\tilde{c}}{\tilde{d}}=\frac{\widetilde{\Phi}_{21}(\Lambda,0;z)}{\widetilde{\Phi}_{22}(\Lambda,0;z)}.$
Step 1: ODE in the $u_{\pm}$ basis. Using $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}u_{+}=-iu_{+}$, $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}u_{-}=iu_{-}$ and $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}\operatorname{Re}p&\operatorname{Im}p\\ \operatorname{Im}p&-\operatorname{Re}p\end{pmatrix}\,u_{+}=ip\,u_{-}$, $\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}\operatorname{Re}p&\operatorname{Im}p\\ \operatorname{Im}p&-\operatorname{Re}p\end{pmatrix}\,u_{-}=-i\overline{p}\,u_{+}$, the propagator satisfies
$\partial_{s}\widetilde{\Phi}_{H}(s,0;z)=\widetilde{M}_{p}(s)\,\widetilde{\Phi}_{H}(s,0;z),\qquad\widetilde{\Phi}_{H}(0,0;z)=I,$
with
$\widetilde{M}_{p}(s)=\begin{pmatrix}\frac{iz}{2}&iz\,\overline{p(s)}\\ -iz\,p(s)&-\frac{iz}{2}\end{pmatrix}$.
Step 2: Duhamel (variation of constants) for traceless perturbations. A perturbation $\Delta H=\begin{pmatrix}\operatorname{Re}q&\operatorname{Im}q\\ \operatorname{Im}q&-\operatorname{Re}q\end{pmatrix}$ corresponds to $p\mapsto p+\varepsilon q$, hence
$\delta\widetilde{M}(s)=\begin{pmatrix}0&iz\,\overline{q(s)}\\ -iz\,q(s)&0\end{pmatrix}$.
Variation of constants gives
$D\widetilde{\Phi}_{H}(\Lambda,0;z)[q]=\int_{0}^{\Lambda}\widetilde{\Phi}_{H}(\Lambda,s;z)\,\delta\widetilde{M}(s)\,\widetilde{\Phi}_{H}(s,0;z)\,ds.$
Step 3: Differentiate the ratio and identify kernels. Write $c:=\widetilde{\Phi}_{21}(\Lambda,0;z)$ and $d:=\widetilde{\Phi}_{22}(\Lambda,0;z)$, so $v=c/d$. Then
$Dv[q]=\frac{Dc[q]-v\,Dd[q]}{d}.$
From the Duhamel formula, taking $(2,1)$ and $(2,2)$ entries yields
$Dc[q]=\int_{0}^{\Lambda}\Bigl(-iz\,q(s)\,\widetilde{\Phi}_{22}(\Lambda,s)\widetilde{\Phi}_{11}(s,0)\;+\;iz\,\overline{q(s)}\,\widetilde{\Phi}_{21}(\Lambda,s)\widetilde{\Phi}_{21}(s,0)\Bigr)\,ds,$
$Dd[q]=\int_{0}^{\Lambda}\Bigl(-iz\,q(s)\,\widetilde{\Phi}_{22}(\Lambda,s)\widetilde{\Phi}_{12}(s,0)\;+\;iz\,\overline{q(s)}\,\widetilde{\Phi}_{21}(\Lambda,s)\widetilde{\Phi}_{22}(s,0)\Bigr)\,ds.$
Substituting into (DcvDd)/d(Dc-vDd)/d and grouping the $q$ and $\overline{q}$ terms gives (35) with
$K^{(1)}_{H}(s;z)$ $:=\frac{\widetilde{\Phi}_{22}(\Lambda,s)\Bigl(\widetilde{\Phi}_{11}(s,0)-v\,\widetilde{\Phi}_{12}(s,0)\Bigr)}{\widetilde{\Phi}_{22}(\Lambda,0)},$ (36)
$K^{(2)}_{H}(s;z)$ $:=-\frac{\widetilde{\Phi}_{21}(\Lambda,s)\Bigl(\widetilde{\Phi}_{21}(s,0)-v\,\widetilde{\Phi}_{22}(s,0)\Bigr)}{\widetilde{\Phi}_{22}(\Lambda,0)}.$ (37)
Step 4: Bounds near $p\equiv 0$. When $p\equiv 0$ one has $\widetilde{\Phi}_{H_{0}}(s,0;z)=\mathrm{diag}(e^{izs/2},e^{-izs/2})$, hence $v_{H_{0},\Lambda}(z)=0$, $K^{(1)}_{H_{0}}(s;z)=e^{izs}$ and $K^{(2)}_{H_{0}}\equiv 0$.
Assume $\|p\|_{L^{1}}\leq\delta$. Writing the ODE in Volterra form,
$\widetilde{\Phi}_{H}(s,t)=\widetilde{\Phi}_{H_{0}}(s,t)+\int_{t}^{s}\widetilde{\Phi}_{H_{0}}(s,r)\,\bigl(\widetilde{M}_{p}(r)-\widetilde{M}_{0}\bigr)\,\widetilde{\Phi}_{H}(r,t)\,dr,$
a Grönwall estimate yields
$\sup_{0\leq t\leq s\leq\Lambda}\|\widetilde{\Phi}_{H}(s,t;z)-\widetilde{\Phi}_{H_{0}}(s,t;z)\|_{\mathrm{op}}\leq C\,\|p\|_{L^{1}(0,\Lambda)},$
with $C=C(z,\Lambda,\varepsilon)$. Choosing $\delta$ so that $C\delta\leq\frac{1}{2}|e^{-iz\Lambda/2}|$ ensures $|\widetilde{\Phi}_{22}(\Lambda,0)|\geq\tfrac{1}{2}|e^{-iz\Lambda/2}|$.
For $K^{(1)}$, the perturbation bound implies uniformly in $s\in[0,\Lambda]$,
$\widetilde{\Phi}_{22}(\Lambda,s)$ $=e^{-iz(\Lambda-s)/2}+O(\|p\|_{L^{1}}),$ (38)
$\widetilde{\Phi}_{11}(s,0)$ $=e^{izs/2}+O(\|p\|_{L^{1}}),$ (39)
$\widetilde{\Phi}_{12}(s,0)$ $=O(\|p\|_{L^{1}}),$ (40)
$v$ $=O(\|p\|_{L^{1}}).$ (41)
hence the numerator of $K^{(1)}$ equals $e^{-iz\Lambda/2}e^{izs}+O(\|p\|_{L^{1}})$ and dividing by $|\widetilde{\Phi}_{22}(\Lambda,0)|\gtrsim e^{\eta\Lambda/2}$ yields $|K^{(1)}_{H}(s;z)|\leq C_{1}|e^{izs}|=C_{1}e^{-\eta s}$.
For $K^{(2)}$, note that at $p=0$ both $\widetilde{\Phi}_{21}(\Lambda,s)$ and $\widetilde{\Phi}_{21}(s,0)$ vanish. The same control implies
$\sup_{s}\bigl|\widetilde{\Phi}_{21}(\Lambda,s)\bigr|$ $\leq C\|p\|_{L^{1}},$ (42)
$\sup_{s}\bigl|\widetilde{\Phi}_{21}(s,0)\bigr|$ $\leq C\|p\|_{L^{1}},$ (43)
$|v|$ $\leq C\|p\|_{L^{1}},$ (44)
$\sup_{s}\bigl|\widetilde{\Phi}_{22}(s,0)\bigr|$ $\leq C.$ (45)
so the numerator of $K^{(2)}$ is $O(\|p\|_{L^{1}}^{2})$ uniformly in $s$, and dividing by $|\widetilde{\Phi}_{22}(\Lambda,0)|\gtrsim e^{\eta\Lambda/2}$ yields $\|K^{(2)}_{H}\|_{\infty}\leq C_{2}\|p\|_{L^{1}}^{2}$.
Finally, if $\operatorname{supp}q\subset[s_{\star},\Lambda]$, then by (35)
$|Dv[q]|\leq|z|\Bigl(\sup_{s\geq s_{\star}}|K^{(1)}_{H}(s;z)|+\|K^{(2)}_{H}\|_{\infty}\Bigr)\|q\|_{L^{1}}\leq|z|\Bigl(C_{1}e^{-\eta s_{\star}}+C_{2}\|p\|_{L^{1}}^{2}\Bigr)\|q\|_{L^{1}}.$
4. Finite-dimensional tangent models: explicit Jacobian and a well-posed minimax depth barrier
4.1. General finite-dimensional tangent parametrizations
Fix $N\geq 1$ and choose basis functions $\{\varphi_{j}\}_{j=1}^{N}\subset L^{1}(0,\Lambda;\mathbb{C})$. Consider the $N$-parameter tangent model
$q_{\theta}(s):=\sum_{j=1}^{N}\theta_{j}\,\varphi_{j}(s),\qquad\theta\in\mathbb{C}^{N},$ (46)
and define the linearized seam operator at $H_{0}$ by
$(T_{x}\theta)_{k}:=Dv_{H_{0},\Lambda}(z_{k})[q_{\theta}],\qquad z_{k}=x_{k}+i\eta.$
Define the Laplace/Fourier features
$\widehat{\varphi}_{j}(z):=\int_{0}^{\Lambda}\varphi_{j}(s)e^{izs}\,ds.$ (47)
Theorem 4.1 (Explicit Jacobian matrix for the general tangent model).
For any nodes $z_{k}=x_{k}+i\eta$,
$(T_{x}\theta)_{k}=-iz_{k}\sum_{j=1}^{N}\theta_{j}\,\widehat{\varphi}_{j}(z_{k}),\qquad 1\leq k\leq M.$ (48)
Equivalently $T_{x}=D_{z}\,\widehat{\Phi}_{x}$ where $D_{z}=\operatorname{diag}(-iz_{k})$ and $(\widehat{\Phi}_{x})_{kj}=\widehat{\varphi}_{j}(z_{k})$.
Proof.
Insert Equation˜46 into Equation˜21 and exchange the finite sum and the integral. ∎
4.2. A well-posed design-minimax bound (uniform in real nodes)
A purely $L^{1}$ assumption does not control $\sup_{x\in\mathbb{R}}|x\,\widehat{\varphi}(x+i\eta)|$ in general, so minimax over unrestricted $x\in\mathbb{R}^{M}$ is not well-posed without additional structure. A mild condition giving uniform control is $\varphi\in W^{1,1}$ (integration by parts).
Assumption 4.2 (Well-posed general-basis model).
Fix $\eta>0$. Assume each $\varphi_{j}$ is absolutely continuous on $0,\Lambda]$ with $\varphi_{j}^{\prime}\in L^{1}(0,\Lambda)$, i.e. $\varphi_{j}\in W^{1,1}(0,\Lambda)$.
Lemma 4.3 (Uniform x-control via integration by parts).
Assume ˜4.2 and fix $z=x+i\eta\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. Then for each $\varphi\in W^{1,1}(0,\Lambda)$,
$z\int_{0}^{\Lambda}\varphi(s)e^{izs}\,ds=-i\Bigl[\varphi(s)e^{izs}\Bigr]_{s=0}^{s=\Lambda}+i\int_{0}^{\Lambda}\varphi^{\prime}(s)e^{izs}\,ds.$ (49)
Consequently,
$\left|z\,\widehat{\varphi}(z)\right|\leq\left|\varphi(\Lambda)\right|e^{-\eta\Lambda}+\left|\varphi(0)\right|+\|\varphi^{\prime}\|_{L^{1}(0,\Lambda)}.$ (50)
Moreover, fix $s_{\star}\in(0,\Lambda]$. If $\varphi$ is supported in $s_{\star},\Lambda]$ and $\varphi^{\prime}\in L^{1}(s_{\star},\Lambda)$, then
$\left|z\,\widehat{\varphi}(z)\right|\leq e^{-\eta s_{\star}}\Bigl(\left|\varphi(s_{\star})\right|+\left|\varphi(\Lambda)\right|+\|\varphi^{\prime}\|_{L^{1}(s_{\star},\Lambda)}\Bigr).$ (51)
Proof.
Since $(e^{izs})^{\prime}=iz\,e^{izs}$, integration by parts yields
$z\int_{0}^{\Lambda}\varphi(s)e^{izs}\,ds=-i\Bigl[\varphi(s)e^{izs}\Bigr]_{s=0}^{s=\Lambda}+i\int_{0}^{\Lambda}\varphi^{\prime}(s)e^{izs}\,ds,$
This is Equation˜49. Taking absolute values and using $|e^{izs}|=e^{-\eta s}\leq 1$, we obtain Equation˜50.
If $\varphi$ is supported in $s_{\star},\Lambda]$, the same computation on $s_{\star},\Lambda]$ together with $\sup_{s\in[s_{\star},\Lambda]}|e^{izs}|\leq e^{-\eta s_{\star}}$ yields Equation˜51. ∎
Definition 4.4 (Design-minimax conditioning (E-optimality)).
For a fixed tangent model (basis functions), define
$\Sigma^{*}_{M}:=\sup_{x\in\mathbb{R}^{M}}\sigma_{\min}(T_{x}),$
with the convention $\sigma_{\min}(T_{x})=0$ if $\operatorname{rank}(T_{x})<N$. Maximizing $\sigma_{\min}(T_{x})$ is the classical E-optimality criterion [17].
Theorem 4.5 (General-basis minimax depth upper bound (well-posed form)).
Fix $\eta>0$ and assume ˜4.2. Let $q_{\theta}$ be the tangent model Equation˜46 and let $T_{x}$ be defined by Equation˜48. Assume there exists an index $j_{\star}$ such that $\varphi_{j_{\star}}$ is supported in $s_{\star},\Lambda]$ with $s_{\star}\in(0,\Lambda]$, and is absolutely continuous on $s_{\star},\Lambda]$ with $\varphi_{j_{\star}}^{\prime}\in L^{1}(s_{\star},\Lambda)$. Then for every design $x\in\mathbb{R}^{M}$,
$\sigma_{\min}(T_{x})\ \leq\ \sqrt{M}\,e^{-\eta s_{\star}}\Bigl(\left|\varphi_{j_{\star}}(s_{\star})\right|+\left|\varphi_{j_{\star}}(\Lambda)\right|+\|\varphi_{j_{\star}}^{\prime}\|_{L^{1}(s_{\star},\Lambda)}\Bigr).$ (52)
In particular, if $s_{\star}$ is the maximal depth start among the basis functions, then
$\Sigma^{*}_{M}\leq\sqrt{M}\,e^{-\eta s_{\star}}\Bigl(\left|\varphi_{j_{\star}}(s_{\star})\right|+\left|\varphi_{j_{\star}}(\Lambda)\right|+\|\varphi_{j_{\star}}^{\prime}\|_{L^{1}(s_{\star},\Lambda)}\Bigr).$
Proof.
Let $e_{j_{\star}}\in\mathbb{C}^{N}$ be the coordinate vector. By the characterization of $\sigma_{\min}$,
$\sigma_{\min}(T_{x})\leq\|T_{x}e_{j_{\star}}\|_{2}.$
Using Equation˜48,
$\left|(T_{x}e_{j_{\star}})_{k}\right|=\left|z_{k}\,\widehat{\varphi}_{j_{\star}}(z_{k})\right|.$
Apply Lemma˜4.3 in the depth-supported form Equation˜51 to bound $\left|z_{k}\widehat{\varphi}_{j_{\star}}(z_{k})\right|$ uniformly over all real $x_{k}$, and then take the $\ell^{2}$ norm over $k=1,\dots,M$ to obtain Equation˜52. ∎
5. A sharp solvable case: block model factorization and design sandwiches
5.1. Block model as a special case
Fix $N\geq 1$ and set $\ell:=\Lambda/N$. Partition $0,\Lambda]$ into $N$ blocks $I_{j}=[j\ell,(j+1)\ell)$, $j=0,\dots,N-1$. A block Hamiltonian is $H(s)\equiv H_{j}$ on $I_{j}$, with $H_{j}\succeq 0$ and $\operatorname{tr}H_{j}=1$. Write
$H_{j}=\frac{1}{2}I+h_{j}\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}+k_{j}\begin{pmatrix}0&1\\[2.0pt] 1&0\end{pmatrix}=\begin{pmatrix}\frac{1}{2}+h_{j}&k_{j}\\ k_{j}&\frac{1}{2}-h_{j}\end{pmatrix},\qquad(h_{j},k_{j})\in\mathbb{R}^{2},$ (53)
and bundle $(h_{j},k_{j})$ into $q_{j}:=h_{j}+ik_{j}\in\mathbb{C}$. (Thus $q_{j}$ parametrizes real-symmetric traceless perturbations via $\begin{pmatrix}\operatorname{Re}q_{j}&\operatorname{Im}q_{j}\\ \operatorname{Im}q_{j}&-\operatorname{Re}q_{j}\end{pmatrix}=h_{j}\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}+k_{j}\begin{pmatrix}0&1\\[2.0pt] 1&0\end{pmatrix}$.)
Theorem 5.1 (Exact block linearization at the free base).
Let $z=x+i\eta\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and let $\Delta q\in\mathbb{C}^{N}$ encode a block perturbation. Then
$Dv_{H_{0},\Lambda}(z)[\Delta q]=-2i\,\sin\!\Bigl(\frac{z\ell}{2}\Bigr)\,\sum_{j=0}^{N-1}\Delta q_{j}\,\exp\!\bigl(iz(j+\tfrac{1}{2})\ell\bigr).$ (54)
Consequently, the linearized seam map $T_{x}:=D\mathcal{S}(0):\mathbb{C}^{N}\to\mathbb{C}^{M}$ satisfies
$(T_{x}\Delta q)_{k}$ $=\gamma_{k}(x_{k})\,\sum_{j=0}^{N-1}\Delta q_{j}\,e^{ix_{k}j\ell}\,w_{j},$ (55)
$\gamma_{k}(x_{k})$ $=-2i\,\sin\!\Bigl(\frac{(x_{k}+i\eta)\ell}{2}\Bigr)\,e^{ix_{k}\ell/2},\qquad w_{j}=e^{-\eta(j+1/2)\ell}.$
Proof.
Specialize Equation˜21 to piecewise-constant $q(s)\equiv\Delta q_{j}$ on $I_{j}$:
$Dv_{H_{0},\Lambda}(z)[\Delta q]=-iz\sum_{j=0}^{N-1}\Delta q_{j}\int_{j\ell}^{(j+1)\ell}e^{izs}\,ds=-iz\sum_{j=0}^{N-1}\Delta q_{j}\,\frac{e^{iz(j+1)\ell}-e^{izj\ell}}{iz}.$
Factor $e^{iz(j+\frac{1}{2})\ell}$ and use $e^{ia}-e^{-ia}=2i\sin a$ to get Equation˜54. Writing $e^{iz(j+1/2)\ell}=e^{ix(j+1/2)\ell}e^{-\eta(j+1/2)\ell}$ yields Equation˜55. ∎
6. Design bounds at the base: minimax sandwiches and an optimal equispaced shift
6.1. A singular-value lemma for left/right factors
We use the following standard singular-value inequality.
Lemma 6.1 (Minimum singular value under left/right factors).
Let $A\in\mathbb{C}^{m\times m}$, $B\in\mathbb{C}^{m\times n}$, and $C\in\mathbb{C}^{n\times n}$. Interpreting $\sigma_{\min}(B)$ as $\sigma_{n}(B)$ (equivalently, $\inf_{\|x\|_{2}=1}\|Bx\|_{2}$), one has
$\sigma_{\min}(ABC)\ \geq\ \sigma_{\min}(A)\,\sigma_{\min}(B)\,\sigma_{\min}(C).$
Proof.
For any $x\in\mathbb{C}^{n}$ with $\|x\|_{2}=1$,
$\|ABCx\|_{2}\geq\sigma_{\min}(A)\,\|BCx\|_{2}\geq\sigma_{\min}(A)\sigma_{\min}(B)\,\|Cx\|_{2}\geq\sigma_{\min}(A)\sigma_{\min}(B)\sigma_{\min}(C)\,\|x\|_{2}.$
Taking the infimum over $\|x\|_{2}=1$ gives the claim. ∎
6.2. Sampling as optimal kernel frames in a finite Paley–Wiener model space
Fix $\ell>0$ and an integer $N\geq 1$. Define the finite-dimensional model space
$\mathcal{V}_{N}:=\Bigl\{F(z)=\sum_{j=0}^{N-1}c_{j}e^{izj\ell}:\ c\in\mathbb{C}^{N}\Bigr\},\qquad\langle F,G\rangle_{\mathcal{V}_{N}}:=\sum_{j=0}^{N-1}c_{j}\overline{d_{j}},$
where $G(z)=\sum_{j=0}^{N-1}d_{j}e^{izj\ell}$. Then $\mathcal{V}_{N}$ is an RKHS on $\mathbb{C}$ with reproducing kernel
$K_{N}(z,w):=\sum_{j=0}^{N-1}e^{izj\ell}\,e^{-i\overline{w}\,j\ell},$
since for $F(z)=\sum_{j=0}^{N-1}c_{j}e^{izj\ell}$ one has
$\langle F,K_{N}(\cdot,w)\rangle_{\mathcal{V}_{N}}=\sum_{j=0}^{N-1}c_{j}\,\overline{e^{-i\overline{w}\,j\ell}}=\sum_{j=0}^{N-1}c_{j}e^{iwj\ell}=F(w).$
Given sampling points $w_{1},\dots,w_{M}\in\mathbb{C}$, define the evaluation operator
$\mathrm{Ev}_{w}:\mathcal{V}_{N}\to\mathbb{C}^{M},\qquad(\mathrm{Ev}_{w}F)_{k}:=F(w_{k}).$
In the coefficient basis $\{e^{izj\ell}\}_{j=0}^{N-1}$, $\mathrm{Ev}_{w}$ is represented by the matrix
[…]
For sampling on a horizontal line $w_{k}=x_{k}+i\eta$ one has the factorization $\mathrm{Ev}_{w}=F_{x}D_{\eta}$ with $(F_{x})_{kj}=e^{ix_{k}j\ell}$ and $D_{\eta}=\operatorname{diag}(e^{-\eta j\ell})$. Thus stability in this finite model is exactly a kernel-frame question for $\{K_{N}(\cdot,w_{k})\}$, governed by $\sigma_{\min}(F_{x})$ and the diagonal depth weights; see Proposition˜9.3. In particular, tight-frame Fourier blocks maximize the lower frame bound (Lemma 6.3), equispaced designs yield tight frames (Lemma 6.10), and the optimal equispaced shift is explicit (Theorem 6.15).
6.3. Fourier block and row-factor bounds
Define the Fourier sampling matrix
$(F_{x})_{kj}:=e^{ix_{k}j\ell},\qquad 1\leq k\leq M,\ \ 0\leq j\leq N-1,$
so that $T_{x}=D_{\gamma}(x)\,F_{x}\,D_{w}$ with $D_{\gamma}(x)=\operatorname{diag}(\gamma_{k}(x_{k}))$ and $D_{w}=\operatorname{diag}(w_{j})$.
Lemma 6.2 (Sine modulus bounds).
For $a\in\mathbb{R}$ and $b>0$,
$|\sin(a+ib)|^{2}=\sin^{2}a+\sinh^{2}b,\qquad\sinh b\leq|\sin(a+ib)|\leq\cosh b.$
Proof.
Use $\sin(a+ib)=\sin a\cosh b+i\cos a\sinh b$ and take moduli. ∎
Lemma 6.3 (Frobenius/tight-frame bound).
For every $x\in\mathbb{R}^{M}$ (with $M\geq N$),
$\sigma_{\min}(F_{x})\leq\sqrt{M}.$
Moreover, $\sigma_{\min}(F_{x})=\sqrt{M}$ holds if and only if $F_{x}^{*}F_{x}=MI_{N}$ (tight-frame columns) [4].
Proof.
$\|F_{x}\|_{F}^{2}=\sum_{k,j}|(F_{x})_{kj}|^{2}=MN$. Let $\sigma_{1}\geq\cdots\geq\sigma_{N}\geq 0$ denote the $N$ singular values of the map $F_{x}:\mathbb{C}^{N}\to\mathbb{C}^{M}$. Then $\sum_{r=1}^{N}\sigma_{r}^{2}=\|F_{x}\|_{F}^{2}=MN$, so $\sigma_{N}^{2}\leq\frac{1}{N}\sum_{r}\sigma_{r}^{2}=M$, i.e. $\sigma_{\min}(F_{x})\leq\sqrt{M}$. Equality forces $\sigma_{1}=\cdots=\sigma_{N}=\sqrt{M}$, which is equivalent to $F_{x}^{*}F_{x}=MI_{N}$. ∎
Theorem 6.4 (Tight frames maximize the lower frame bound).
Assume $M\geq N$. Then $\sigma_{\min}(F_{x})\leq\sqrt{M}$ for every $x\in\mathbb{R}^{M}$, with equality if and only if $F_{x}^{*}F_{x}=MI_{N}$. Equivalently, the kernels $\{K_{N}(\cdot,x_{k})\}_{k=1}^{M}$ form a tight frame for $\mathcal{V}_{N}$.
Proof.
This is exactly Lemma 6.3, together with the identification of $F_{x}$ as an evaluation matrix for $\mathcal{V}_{N}$ from Section˜6.2. ∎
6.4. A rigidity statement for the tight-frame Fourier block (square case)
Theorem 6.5 (Rigidity for the Fourier block in the square case).
Assume $M=N$. Then $\sigma_{\min}(F_{x})\leq\sqrt{N}$. Moreover, $\sigma_{\min}(F_{x})=\sqrt{N}$ if and only if, after reordering,
$x_{k}\ell\equiv\beta+\frac{2\pi(k-1)}{N}\pmod{2\pi}$
for some $\beta\in\mathbb{R}$, i.e. $\{e^{ix_{k}\ell}\}_{k=1}^{N}$ is the set of $N$th roots of unity up to a common phase.
Proof.
We have $\|F_{x}\|_{F}^{2}=\sum_{k=1}^{N}\sum_{j=0}^{N-1}|e^{ix_{k}j\ell}|^{2}=N^{2}$. Let $\sigma_{1}\geq\cdots\geq\sigma_{N}\geq 0$ be the singular values of $F_{x}$. Then $\sum_{r=1}^{N}\sigma_{r}^{2}=\|F_{x}\|_{F}^{2}=N^{2}$, hence $\sigma_{\min}^{2}\leq N$ and $\sigma_{\min}\leq\sqrt{N}$.
If $\sigma_{\min}(F_{x})=\sqrt{N}$, then all singular values equal $\sqrt{N}$, so $F_{x}^{*}F_{x}=NI_{N}$. Let $\zeta_{k}:=e^{ix_{k}\ell}$. The $(j,j^{\prime})$ entry of $F_{x}^{*}F_{x}$ equals $\sum_{k=1}^{N}\zeta_{k}^{j^{\prime}-j}$, so $F_{x}^{*}F_{x}=NI_{N}$ is equivalent to
$\sum_{k=1}^{N}\zeta_{k}^{r}=0,\qquad r=1,2,\dots,N-1.$
Let $p(z):=\prod_{k=1}^{N}(z-\zeta_{k})=z^{N}-e_{1}z^{N-1}+e_{2}z^{N-2}-\cdots+(-1)^{N}e_{N}$, where $e_{j}$ are the elementary symmetric polynomials in $\zeta_{1},\dots,\zeta_{N}$ and $e_{0}:=1$. Set $s_{r}:=\sum_{k=1}^{N}\zeta_{k}^{r}$ (power sums). Newton’s identities give, for $1\leq m\leq N$,
$m\,e_{m}=\sum_{r=1}^{m}(-1)^{r-1}e_{m-r}\,s_{r}.$
For $m\leq N-1$ we have $s_{1}=\cdots=s_{m}=0$, hence $e_{m}=0$ (inductively on $m$). Therefore $e_{1}=\cdots=e_{N-1}=0$ and
$p(z)=z^{N}+(-1)^{N}e_{N}=z^{N}-a,\qquad a:=(-1)^{N+1}e_{N}.$
Because $|\zeta_{k}|=1$, we have $|e_{N}|=|\prod_{k}\zeta_{k}|=1$, hence $|a|=1$. Thus $\{\zeta_{k}\}$ is exactly the set of roots of $z^{N}=a$, i.e. $\zeta_{k}=e^{i\beta/N}\omega^{k-1}$ with $\omega=e^{2\pi i/N}$ up to permutation, which is the stated congruence.
Conversely, if $\{\zeta_{k}\}$ is a rotated set of $N$th roots of unity, then $F_{x}$ is a DFT matrix up to diagonal phases, so $F_{x}^{*}F_{x}=NI_{N}$ and $\sigma_{\min}(F_{x})=\sqrt{N}$. ∎
Theorem 6.6 (Fourier coefficients control near-tightness of Fx).
Let $M\geq N$ and set $\theta_{k}:=x_{k}\ell\ (\mathrm{mod}\ 2\pi)$. Let $\mu:=\frac{1}{M}\sum_{k=1}^{M}\delta_{\theta_{k}}$ and $\widehat{\mu}(r):=\frac{1}{M}\sum_{k=1}^{M}e^{ir\theta_{k}}$. Define the Gram matrix
$A:=\frac{1}{M}F_{x}^{*}F_{x}\in\mathbb{C}^{N\times N}.$
Then
$A_{jj^{\prime}}=\widehat{\mu}(j^{\prime}-j),\qquad 0\leq j,j^{\prime}\leq N-1,$
so $(A-I)_{jj^{\prime}}=\widehat{\mu}(j^{\prime}-j)$ for $j\neq j^{\prime}$ and $(A-I)_{jj}=0$. Moreover,
$\sup_{1\leq r\leq N-1}|\widehat{\mu}(r)|\ \leq\ \|A-I\|_{2\to 2}\ \leq\ \|A-I\|_{1\to 1}\ \leq\ 2\sum_{r=1}^{N-1}|\widehat{\mu}(r)|.$ (56)
Consequently:
  1. (i)
    If $\|A-I\|_{2\to 2}\leq\varepsilon$, then $|\widehat{\mu}(r)|\leq\varepsilon$ for all $1\leq r\leq N-1$.
  2. (ii)
    If $\sum_{r=1}^{N-1}|\widehat{\mu}(r)|\leq\varepsilon/2$ with $\varepsilon\in(0,1)$, then $\|A-I\|_{2\to 2}\leq\varepsilon$ and hence $\sigma_{\min}(F_{x})\geq\sqrt{M}\,(1-\varepsilon)^{1/2}$.
Proof.
Compute
$A_{jj^{\prime}}=\frac{1}{M}\sum_{k=1}^{M}e^{-ix_{k}j\ell}e^{ix_{k}j^{\prime}\ell}=\frac{1}{M}\sum_{k=1}^{M}e^{i(j^{\prime}-j)\theta_{k}}=\widehat{\mu}(j^{\prime}-j).$
For the first inequality in (56), fix $j\neq j^{\prime}$ and note $|(A-I)_{jj^{\prime}}|=|e_{j}^{*}(A-I)e_{j^{\prime}}|\leq\|A-I\|_{2\to 2}$; taking the supremum over $r=j^{\prime}-j$ yields $\sup_{1\leq r\leq N-1}|\widehat{\mu}(r)|\leq\|A-I\|_{2\to 2}$.
The inequality $\|B\|_{2\to 2}\leq\|B\|_{1\to 1}$ holds for all matrices. For the last inequality, recall $\|A-I\|_{1\to 1}=\max_{j^{\prime}}\sum_{j}|(A-I)_{jj^{\prime}}|$. Fix a column $j^{\prime}$. The diagonal term is 0 and all off-diagonal entries are $\widehat{\mu}(r)$ with $r\neq 0$. Thus
$\sum_{j}|(A-I)_{jj^{\prime}}|\leq\sum_{r\neq 0,\ |r|\leq N-1}|\widehat{\mu}(r)|=2\sum_{r=1}^{N-1}|\widehat{\mu}(r)|,$
using $|\widehat{\mu}(-r)|=|\overline{\widehat{\mu}(r)}|=|\widehat{\mu}(r)|$.
For (ii), $\|A-I\|_{2\to 2}\leq\varepsilon$ implies $\lambda_{\min}(A)\geq 1-\varepsilon$ since $A$ is Hermitian. Since $A=\frac{1}{M}F_{x}^{*}F_{x}$, we obtain $\sigma_{\min}(F_{x})^{2}/M=\lambda_{\min}(A)\geq 1-\varepsilon$. ∎
6.5. Universal depth-barrier upper bound (block model)
Theorem 6.7 (Universal design upper bound and depth barrier (block model)).
Fix $\eta>0$ and integers $M\geq N\geq 1$. For any design $x\in\mathbb{R}^{M}$,
$\sigma_{\min}(T_{x})\ \leq\ 2\sqrt{M}\,\cosh\!\Bigl(\frac{\eta\ell}{2}\Bigr)\,e^{-\eta(\Lambda-\ell/2)}.$ (57)
Consequently,
$\Sigma^{*}_{M}(\eta;\Lambda,N)\ \leq\ 2\sqrt{M}\,\cosh\!\Bigl(\frac{\eta\ell}{2}\Bigr)\,e^{-\eta(\Lambda-\ell/2)}.$
Proof.
Let $e_{N-1}\in\mathbb{C}^{N}$ be supported on the deepest block. Then $\sigma_{\min}(T_{x})\leq\|T_{x}e_{N-1}\|_{2}$. From Equation˜55, $(T_{x}e_{N-1})_{k}=\gamma_{k}(x_{k})e^{ix_{k}(N-1)\ell}w_{N-1}$, hence
$\|T_{x}e_{N-1}\|_{2}^{2}=w_{N-1}^{2}\sum_{k=1}^{M}|\gamma_{k}(x_{k})|^{2}\leq w_{N-1}^{2}\,M\Bigl(\max_{k}|\gamma_{k}(x_{k})|\Bigr)^{2}.$
Now $|\gamma_{k}(x_{k})|=2|\sin((x_{k}+i\eta)\ell/2)|\leq 2\cosh(\eta\ell/2)$ by Lemma˜6.2, and $w_{N-1}=e^{-\eta(\Lambda-\ell/2)}$. Taking square-roots yields Equation˜57. ∎
Corollary 6.8 (Near-saturation of the universal upper bound forces near-sine-maximization).
Assume the square block model $M=N$ and fix $\eta,\Lambda,N$. Let $T_{x}=D_{\gamma}(x)\,F_{x}\,D_{w}$ be the Jacobian at $H_{0}$ and set $b:=\eta\ell/2$. Assume the universal pointwise bound $|\gamma_{k}(x_{k})|\leq 2\cosh(b)$ holds for all $k$, and assume
$\sigma_{\min}(T_{x})\ \geq\ (1-\varepsilon)\,2\sqrt{N}\cosh(b)\,e^{-\eta(\Lambda-\ell/2)}.$
Then
$\frac{1}{N}\sum_{k=1}^{N}\Bigl(\frac{|\gamma_{k}(x_{k})|}{2\cosh(b)}\Bigr)^{2}\ \geq\ (1-\varepsilon)^{2}.$
In particular, for every $\tau\in(0,1)$,
$\Bigl\{k:\ |\gamma_{k}(x_{k})|\leq 2\cosh(b)\sqrt{1-\tau}\Bigr\}\ \leq\ \frac{\varepsilon(2-\varepsilon)}{\tau}\,N.$
Proof.
Let $e_{N-1}$ be the deepest-coordinate unit vector in $\mathbb{C}^{N}$. Then $\sigma_{\min}(T_{x})\leq\|T_{x}e_{N-1}\|_{2}$. Since $T_{x}=D_{\gamma}F_{x}D_{w}$ and $F_{x}e_{N-1}=(e^{ix_{k}(N-1)\ell})_{k=1}^{N}$, we have
$T_{x}e_{N-1}=w_{N-1}\,D_{\gamma}\,(e^{ix_{k}(N-1)\ell})_{k=1}^{N},$
hence
$\|T_{x}e_{N-1}\|_{2}^{2}=w_{N-1}^{2}\sum_{k=1}^{N}|\gamma_{k}(x_{k})|^{2}.$
Therefore
$\sigma_{\min}(T_{x})^{2}\leq w_{N-1}^{2}\sum_{k=1}^{N}|\gamma_{k}(x_{k})|^{2}.$
Using the hypothesis and $w_{N-1}=e^{-\eta(\Lambda-\ell/2)}$ yields
$(1-\varepsilon)^{2}\,4N\cosh^{2}(b)\,w_{N-1}^{2}\leq w_{N-1}^{2}\sum_{k=1}^{N}|\gamma_{k}(x_{k})|^{2},$
and canceling $w_{N-1}^{2}$ gives the average bound.
For the counting statement, set $a_{k}:=\bigl(|\gamma_{k}|/(2\cosh b)\bigr)^{2}\in[0,1]$. If $a_{k}\leq 1-\tau$ on a set $S$, then
$\frac{1}{N}\sum_{k=1}^{N}a_{k}\leq\frac{|S|}{N}(1-\tau)+\Bigl(1-\frac{|S|}{N}\Bigr)\cdot 1=1-\tau\frac{|S|}{N}.$
Combine with $\frac{1}{N}\sum_{k}a_{k}\geq(1-\varepsilon)^{2}=1-\varepsilon(2-\varepsilon)$ to obtain $|S|\leq\frac{\varepsilon(2-\varepsilon)}{\tau}N$. ∎
6.6. Shifted-DFT and an optimal equispaced shift
Definition 6.9 (Equispaced design family).
Fix $M\geq 1$ and define, for a shift $\alpha\in\mathbb{R}$,
$x_{k}(\alpha):=\frac{2\pi(k-1)}{M\ell}+\frac{2\alpha}{\ell},\qquad k=1,\dots,M,\qquad z_{k}(\alpha)=x_{k}(\alpha)+i\eta.$
Equivalently,
$\frac{x_{k}(\alpha)\ell}{2}=\alpha+\frac{\pi(k-1)}{M}.$
Lemma 6.10 (Equispaced designs are tight for the Fourier block).
Assume $N\leq M$. For every $\alpha\in\mathbb{R}$, the Fourier block $F_{x(\alpha)}$ satisfies
$F_{x(\alpha)}^{*}F_{x(\alpha)}=MI_{N},\qquad\text{hence}\qquad\sigma_{\min}(F_{x(\alpha)})=\sqrt{M}.$
Proof.
For $0\leq j,j^{\prime}\leq N-1$,
$(F^{*}F)_{jj^{\prime}}=\sum_{k=1}^{M}e^{-ix_{k}(\alpha)j\ell}e^{ix_{k}(\alpha)j^{\prime}\ell}=\sum_{k=1}^{M}\exp\!\Bigl(\frac{2\pi i(k-1)}{M}(j^{\prime}-j)\Bigr),$
since the $\alpha$-shift cancels. If $j=j^{\prime}$ the sum is $M$. If $j\neq j^{\prime}$, then $1\leq|j^{\prime}-j|\leq N-1\leq M-1$, so the sum of $M$th roots of unity is 0. ∎
Theorem 6.11 (Equispaced sampling gives an optimal tight kernel frame).
Assume $M\geq N$ and consider the equispaced family $x(\alpha)$ from Definition˜6.9. Then $F_{x(\alpha)}^{*}F_{x(\alpha)}=MI_{N}$ and $\sigma_{\min}(F_{x(\alpha)})=\sqrt{M}$.
Proof.
This is Lemma 6.10. ∎
Lemma 6.12 (Worst-case row factor over equispaced shifts).
Fix $b>0$ and integer $M\geq 2$. Define
$\psi_{M}(\alpha):=\min_{1\leq k\leq M}\left|\sin\!\Bigl(\alpha+\frac{\pi(k-1)}{M}+ib\Bigr)\right|.$
Then
$\sup_{\alpha\in\mathbb{R}}\psi_{M}(\alpha)=\sqrt{\sinh^{2}b+\sin^{2}\!\Bigl(\frac{\pi}{2M}\Bigr)},$
and every maximizer $\alpha$ is congruent to $\pi/(2M)$ modulo $\pi/M$.
Proof.
By Lemma˜6.2,
$\left|\sin(a+ib)\right|^{2}=\sinh^{2}b+\sin^{2}a.$
Hence maximizing $\psi_{M}(\alpha)$ is equivalent to maximizing
$m(\alpha):=\min_{1\leq k\leq M}\sin^{2}\!\Bigl(\alpha+\frac{\pi(k-1)}{M}\Bigr).$
Let $G_{\alpha}:=\{\alpha+\pi(k-1)/M:\ k=1,\dots,M\}$ considered modulo $\pi$. This is an $\pi/M$-spaced grid in $\mathbb{R}/\pi\mathbb{Z}$. Let $d(\alpha)$ be the distance in $\mathbb{R}/\pi\mathbb{Z}$ from 0 (a zero of $\sin$) to the nearest point in $G_{\alpha}$. Then $\min_{k}|\sin(\alpha+\pi(k-1)/M)|=\sin(d(\alpha))$ with $d(\alpha)\in[0,\pi/2]$. Since grid spacing is $\pi/M$, the maximal possible nearest-neighbor distance to 0 is $\pi/(2M)$, attained exactly when 0 lies at the midpoint between two consecutive grid points, i.e. when $\alpha\equiv\pi/(2M)\ (\mathrm{mod}\ \pi/M)$. Thus $\sup_{\alpha}m(\alpha)=\sin^{2}(\pi/(2M))$. Substituting back gives the claim. ∎
Theorem 6.13 (Shifted-DFT design: explicit conditioning lower bound).
Fix $\eta>0$ and integers $M\geq N\geq 1$ with $M\geq 2$. Choose the half-shifted equispaced design:
$x_{k}:=\frac{2\pi(k-\tfrac{1}{2})}{M\ell},\quad k=1,\dots,M,\qquad z_{k}=x_{k}+i\eta.$
Then:
  1. (i)
    (Tight-frame Fourier block) $F_{x}^{*}F_{x}=MI_{N}$, hence $\sigma_{\min}(F_{x})=\sqrt{M}$.
  2. (ii)
    (Exact row-factor minimum)
    $\min_{1\leq k\leq M}|\gamma_{k}(x_{k})|=2\sqrt{\sinh^{2}(\eta\ell/2)+\sin^{2}(\pi/(2M))}.$
  3. (iii)
    (Conditioning lower bound)
    $\sigma_{\min}(T_{x})\ \geq\ 2\sqrt{M}\,\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2M}\Bigr)}\,e^{-\eta(\Lambda-\ell/2)}.$ (58)
Proof.
(i) This is Lemma˜6.10 with $\alpha=\pi/(2M)$. (ii) Here $(x_{k}\ell)/2=\pi(k-\tfrac{1}{2})/M$. By Lemma˜6.2,
$|\sin((x_{k}+i\eta)\ell/2)|^{2}=\sin^{2}(\pi(k-\tfrac{1}{2})/M)+\sinh^{2}(\eta\ell/2).$
The minimum over $k$ is attained at $k=1$ or $k=M$, giving $\sin(\pi/(2M))$. (iii) Factor $T_{x}=D_{\gamma}F_{x}D_{w}$ and apply Lemma˜6.1. We have $\sigma_{\min}(F_{x})=\sqrt{M}$ by (i), $\sigma_{\min}(D_{\gamma})=\min_{k}|\gamma_{k}(x_{k})|$, and $\sigma_{\min}(D_{w})=\min_{j}w_{j}=e^{-\eta(\Lambda-\ell/2)}$. Combine with (ii). ∎
Theorem 6.14 (Minimax sandwich (block model)).
Fix $\eta>0$ and integers $M\geq N\geq 1$ with $M\geq 2$. Then
$2\sqrt{M}\,\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2M}\Bigr)}\,e^{-\eta(\Lambda-\ell/2)}$ $\leq\Sigma^{*}_{M}(\eta;\Lambda,N)$
$\leq 2\sqrt{M}\,\cosh\!\Bigl(\frac{\eta\ell}{2}\Bigr)\,e^{-\eta(\Lambda-\ell/2)}.$
In particular, for fixed $\eta,\Lambda$ and $\ell=\Lambda/N$,
$\Sigma^{*}_{M}(\eta;\Lambda,N)\asymp\sqrt{M}\,e^{-\eta(\Lambda-\ell/2)}$
with explicit constants.
Proof.
Lower bound: Theorem˜6.13. Upper bound: Theorem˜6.7. ∎
Theorem 6.15 (Exact shift-optimization among equispaced Fourier-tight designs).
Fix integers $M\geq N\geq 1$ with $M\geq 2$ and $\eta>0$. For the equispaced family $x(\alpha)$ from Definition˜6.9, the Fourier block is tight for every shift $\alpha$, and
$\max_{\alpha\in\mathbb{R}}\ \min_{1\leq k\leq M}\left|\gamma_{k}(x_{k}(\alpha))\right|=2\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2M}\Bigr)}.$
Every maximizer is congruent to $\alpha=\pi/(2M)$ modulo $\pi/M$.
Proof.
This is a direct reformulation of Lemma˜6.12 with $b=\eta\ell/2$ and $\gamma_{k}(x_{k})=-2i\sin(z_{k}\ell/2)e^{ix_{k}\ell/2}$. ∎
7. Quantitative local inversion (square case)
7.1. Block-positive domain
Fix $\varepsilon\in(0,\tfrac{1}{2})$ and define
$U_{\varepsilon}:=\Bigl\{\theta\in\mathbb{R}^{2N}:\ \max_{0\leq j\leq N-1}\sqrt{h_{j}^{2}+k_{j}^{2}}<\frac{1}{2}-\varepsilon\Bigr\}.$
Then each block $H_{j}(\theta)$ has eigenvalues in $(\varepsilon,1-\varepsilon)$.
7.2. First and second variations and explicit bounds
For $\theta\in U_{\varepsilon}$ define $\Phi_{\theta}(s,t;z):=\Phi_{H(\theta)}(s,t;z)$. Let $\Delta\theta\in\mathbb{R}^{2N}$ induce $\Delta H$. On each block $I_{j}$ one has $\Delta H=\Delta h_{j}\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}+\Delta k_{j}\begin{pmatrix}0&1\\[2.0pt] 1&0\end{pmatrix}$.
Lemma 7.1 (Duhamel formulas and global bounds).
Fix $z\in\mathbb{C}$. For each $\theta\in U_{\varepsilon}$ and directions $\Delta\theta,\Delta\theta_{1},\Delta\theta_{2}$:
$D\Phi_{\theta}(\Lambda,0;z)[\Delta\theta]$ $=-\,z\int_{0}^{\Lambda}\Phi_{\theta}(\Lambda,s;z)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H(s)\,\Phi_{\theta}(s,0;z)\,ds,$ (59)
$D^{2}\Phi_{\theta}(\Lambda,0;z)\bigl[\Delta\theta_{1},\Delta\theta_{2}\bigr]$ $=z^{2}\int_{0}^{\Lambda}\int_{0}^{s}\Phi_{\theta}(\Lambda,s)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H_{1}(s)$ (60)
$\qquad\cdot\Phi_{\theta}(s,u)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H_{2}(u)\,\Phi_{\theta}(u,0)\,du\,ds$ (61)
$\quad+\,z^{2}\int_{0}^{\Lambda}\int_{0}^{s}\Phi_{\theta}(\Lambda,s)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H_{2}(s)$ (62)
$\qquad\cdot\Phi_{\theta}(s,u)\,\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Delta H_{1}(u)\,\Phi_{\theta}(u,0)\,du\,ds.$ (63)
Moreover, with respect to the Euclidean norm on $\mathbb{R}^{2N}$ and operator norm on $\mathbb{R}^{2\times 2}$,
$\|D\Phi_{\theta}(\Lambda,0;z)\|_{2\to\mathrm{op}}\leq C_{1}(z,\Lambda),\qquad\|D^{2}\Phi_{\theta}(\Lambda,0;z)\|_{2\to 2\to\mathrm{op}}\leq C_{2}(z,\Lambda),$ (64)
where one may take
$C_{1}(z,\Lambda):=|z|\Lambda\,e^{2|z|\Lambda},\qquad C_{2}(z,\Lambda):=|z|^{2}\Lambda^{2}\,e^{3|z|\Lambda}.$
Proof.
The identities are standard first and second variation-of-constants formulas; see [5]. For the bounds, use Lemma˜2.2:
$\left\|\Phi_{\theta}(\Lambda,s)\right\|_{\mathrm{op}}\leq e^{|z|(\Lambda-s)},\qquad\left\|\Phi_{\theta}(s,0)\right\|_{\mathrm{op}}\leq e^{|z|s}.$
For each $s\in I_{j}$ the perturbation satisfies $\|\Delta H(s)\|_{\mathrm{op}}=\sqrt{(\Delta h_{j})^{2}+(\Delta k_{j})^{2}}\leq\|\Delta\theta\|_{2}$. Thus
$\left\|D\Phi_{\theta}(\Lambda,0;z)[\Delta\theta]\right\|_{\mathrm{op}}\leq|z|\int_{0}^{\Lambda}e^{|z|(\Lambda-s)}\,\|\Delta H(s)\|_{\mathrm{op}}\,e^{|z|s}\,ds\leq|z|\Lambda e^{|z|\Lambda}\,\|\Delta\theta\|_{2}.$
We inflate $e^{|z|\Lambda}$ to $e^{2|z|\Lambda}$ to simplify downstream constants. For $D^{2}\Phi$, each of the two symmetric terms is bounded by
$|z|^{2}\int_{0}^{\Lambda}\!\!\int_{0}^{s}e^{|z|(\Lambda-s)}\|\Delta H_{1}(s)\|_{\mathrm{op}}e^{|z|(s-u)}\|\Delta H_{2}(u)\|_{\mathrm{op}}e^{|z|u}\,du\,ds\leq\frac{1}{2}|z|^{2}\Lambda^{2}e^{|z|\Lambda}\,\|\Delta\theta_{1}\|_{2}\|\Delta\theta_{2}\|_{2},$
Since there are two such symmetric terms, their sum is bounded by $|z|^{2}\Lambda^{2}e^{|z|\Lambda}\|\Delta\theta_{1}\|_{2}\|\Delta\theta_{2}\|_{2}$, and we again inflate the exponential constant to the stated $C_{2}$. ∎
7.3. A denominator margin near the base
Fix nodes $z_{k}=x_{k}+i\eta$ and set $Z:=\max_{1\leq k\leq N}|z_{k}|$. For each $k$, define $d_{k}(\theta):=D_{H(\theta)}(z_{k})-iB_{H(\theta)}(z_{k})$.
Lemma 7.2 (Explicit nonvanishing margin near the base).
Let $r_{d}>0$ satisfy
$r_{d}\ \leq\ \frac{e^{\eta\Lambda/2}}{8\,C_{1}(Z,\Lambda)}.$ (65)
Then for all $\theta$ with $\|\theta\|_{2}\leq r_{d}$ and all $k=1,\dots,N$,
$|d_{k}(\theta)|\ \geq\ \frac{1}{2}e^{\eta\Lambda/2}.$
Proof.
By the mean value theorem and Lemma˜7.1,
$\|\Phi_{\theta}(\Lambda,0;z_{k})-\Phi_{0}(\Lambda,0;z_{k})\|_{\mathrm{op}}\leq C_{1}(Z,\Lambda)\,\|\theta\|_{2}\leq C_{1}(Z,\Lambda)\,r_{d}.$
In $2\times 2$, entrywise changes are bounded by $\|\cdot\|_{\mathrm{op}}$, hence
$|d_{k}(\theta)-d_{k}(0)|\leq|D(\theta)-D(0)|+|B(\theta)-B(0)|\leq 2\|\Phi_{\theta}-\Phi_{0}\|_{\mathrm{op}}\leq 2C_{1}(Z,\Lambda)\,r_{d}.$
Using Equation˜16 gives $|d_{k}(0)|=e^{\eta\Lambda/2}$. Under Equation˜65, $|d_{k}(\theta)-d_{k}(0)|\leq e^{\eta\Lambda/2}/2$, hence $|d_{k}(\theta)|\geq e^{\eta\Lambda/2}/2$. ∎
7.4. Uniform $C^{2}$ bounds for $\mathcal{S}$ near the base
Lemma 7.3 (C2 quotient bounds).
Let $U\subset\mathbb{R}^{d}$ be open, $K\Subset U$ compact, and $n,d\in C^{2}(U;\mathbb{C})$. Assume $\inf_{x\in K}|d(x)|\geq\delta_{0}>0$. Define $m:=n/d$. Then $m\in C^{2}(U;\mathbb{C})$ and for all $x\in K$,
$\|Dm(x)\|\ \leq\ \frac{\|Dn\|_{K}}{\delta_{0}}+\frac{\|n\|_{K}\|Dd\|_{K}}{\delta_{0}^{2}},$
and
$\|D^{2}m(x)\|\ \leq\ \frac{\|D^{2}n\|_{K}}{\delta_{0}}+\frac{2\|Dn\|_{K}\|Dd\|_{K}}{\delta_{0}^{2}}+\frac{\|n\|_{K}\|D^{2}d\|_{K}}{\delta_{0}^{2}}+\frac{2\|n\|_{K}\|Dd\|_{K}^{2}}{\delta_{0}^{3}}.$
Proof.
Write $m=n\cdot d^{-1}$ and use $D(d^{-1})=-d^{-2}Dd$ and $D^{2}(d^{-1})=2d^{-3}(Dd)\otimes(Dd)-d^{-2}D^{2}d$, then apply the product rule and bound each term using $|d|\geq\delta_{0}$ on $K$. ∎
Lemma 7.4 (Uniform bound for second derivatives of v near the base).
Fix $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and let $C_{1},C_{2}$ be as in Lemma˜7.1. Assume $\theta$ lies in a set where $|d(\theta)|\geq\delta_{0}>0$ for $d(\theta)=D_{H(\theta)}(z)-iB_{H(\theta)}(z)$. Then $\theta\mapsto v_{H(\theta),\Lambda}(z)$ is $C^{2}$ there and satisfies
$\|D^{2}v_{H(\theta),\Lambda}(z)\|\ \leq\ 4\|Dm(\theta)\|^{2}+2\|D^{2}m(\theta)\|,$
where $m(\theta)=(iA-C)/(D-iB)$. Moreover, on any compact set $K$ where $|d|\geq\delta_{0}$,
$\sup_{\theta\in K}\|Dm(\theta)\|\leq\frac{2C_{1}}{\delta_{0}}+\frac{(2e^{|z|\Lambda})(2C_{1})}{\delta_{0}^{2}},$
and
$\sup_{\theta\in K}\|D^{2}m(\theta)\|\leq\frac{2C_{2}}{\delta_{0}}+\frac{2(2C_{1})(2C_{1})}{\delta_{0}^{2}}+\frac{(2e^{|z|\Lambda})(2C_{2})}{\delta_{0}^{2}}+\frac{2(2e^{|z|\Lambda})(2C_{1})^{2}}{\delta_{0}^{3}}.$
Proof.
Write $m=n/d$ with n=iACn=iA-C and d=DiBd=D-iB. On any set where $|d|\geq\delta_{0}$, apply Lemma˜7.3. For bounds, note $\|\Phi_{\theta}(\Lambda,0;z)\|_{\mathrm{op}}\leq e^{|z|\Lambda}$ by Lemma˜2.2, hence $|n|\leq 2e^{|z|\Lambda}$. Also in $2\times 2$, entrywise derivatives are bounded by operator-norm derivatives, and Lemma˜7.1 gives $\|Dn\|\leq 2C_{1}$, $\|Dd\|\leq 2C_{1}$, $\|D^{2}n\|\leq 2C_{2}$, $\|D^{2}d\|\leq 2C_{2}$. Finally, $v=\mathcal{C}(m)$ with $\mathcal{C}(w)=(w-i)/(w+i)$. Since $m(\theta)\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, writing $m=x+iy$ with $y>0$ gives $|m+i|^{2}=x^{2}+(y+1)^{2}\geq 1$, hence $|\mathcal{C}^{\prime}(m)|=\frac{2}{|m+i|^{2}}\leq 2$ and $|\mathcal{C}^{\prime\prime}(m)|=\frac{4}{|m+i|^{3}}\leq 4$. The second-derivative chain rule yields $\|D^{2}v\|\leq\sup|\mathcal{C}^{\prime\prime}|\|Dm\|^{2}+\sup|\mathcal{C}^{\prime}|\|D^{2}m\|\leq 4\|Dm\|^{2}+2\|D^{2}m\|$. ∎
Theorem 7.5 (Uniform second-derivative bound for S and Lipschitz constant for DS).
Assume $M=N$ and fix height $\eta>0$ and nodes $z_{k}=x_{k}+i\eta$. Let $Z:=\max_{k}|z_{k}|$ and choose $r_{d}$ as in Lemma˜7.2. Set $\delta_{0}:=\tfrac{1}{2}e^{\eta\Lambda/2}$. Then on the Euclidean ball $B(0,r_{d})\subset\mathbb{R}^{2N}$,
$\sup_{\theta\in B(0,r_{d})}\|D^{2}\mathcal{S}(\theta)\|_{2\to 2\to 2}\ \leq\ \sqrt{N}\,\max_{1\leq k\leq N}\Bigl(\sup_{\theta\in B(0,r_{d})}\|D^{2}v_{H(\theta),\Lambda}(z_{k})\|\Bigr)=:B<\infty,$
and consequently
$\|D\mathcal{S}(\theta)-D\mathcal{S}(\tilde{\theta})\|_{2\to 2}\ \leq\ B\,\|\theta-\tilde{\theta}\|_{2},\qquad\theta,\tilde{\theta}\in B(0,r_{d}).$
Proof.
By Lemma˜7.2, $|d_{k}(\theta)|\geq\delta_{0}$ for $\theta\in B(0,r_{d})$ and all $k$. Apply Lemma˜7.4 componentwise and combine using $\|(a_{k})\|_{2}\leq\sqrt{N}\max_{k}|a_{k}|$. The Lipschitz bound follows from the mean value theorem in Banach spaces. ∎
Lemma 7.6 (Quantitative inverse function theorem in contraction form).
Let $F:B(0,r)\subset\mathbb{R}^{d}\to\mathbb{R}^{d}$ be $C^{1}$ with $F(0)=0$, and let $T:=DF(0)$ be invertible with $M_{0}:=\|T^{-1}\|_{2\to 2}$. Assume there is $B\geq 0$ such that
$\|DF(x)-DF(y)\|_{2\to 2}\leq B\|x-y\|_{2},\qquad x,y\in B(0,r),$
and that
$BM_{0}r\leq\frac{1}{2}.$
Set $\delta:=\frac{r}{2M_{0}}$. Then for every $y\in\mathbb{R}^{d}$ with $\|y\|_{2}\leq\delta$ there exists a unique $x\in B(0,r)$ such that $F(x)=y$. Moreover the local inverse satisfies the Lipschitz bound
$\|F^{-1}(y)-F^{-1}(\tilde{y})\|_{2}\leq 2M_{0}\,\|y-\tilde{y}\|_{2},\qquad\|y\|_{2},\ \|\tilde{y}\|_{2}\leq\delta.$
Proof.
This is a standard contraction-form inverse function theorem argument; see, e.g., [7]. For fixed $y$ define the map $G_{y}(x):=x-T^{-1}(F(x)-y)$ on $B(0,r)$. For $x\in B(0,r)$,
$DG_{y}(x)=I-T^{-1}DF(x)=T^{-1}(T-DF(x)).$
Using $\|DF(x)-T\|_{2\to 2}\leq B\|x\|_{2}\leq Br$, we have
$\|DG_{y}(x)\|_{2\to 2}\leq M_{0}\,\|DF(x)-T\|_{2\to 2}\leq BM_{0}r\leq\frac{1}{2},$
so $G_{y}$ is a contraction on $B(0,r)$. Also, by the integral form of the mean value theorem and the Lipschitz bound on $DF$,
$F(x)-Tx$ $=\int_{0}^{1}(DF(tx)-T)x\,dt,$
$\|F(x)-Tx\|_{2}$ $\leq\int_{0}^{1}Bt\|x\|_{2}^{2}\,dt=\frac{B}{2}\|x\|_{2}^{2}\leq\frac{B}{2}r\|x\|_{2}.$
and hence for $\|x\|_{2}\leq r$ and $\|y\|_{2}\leq\delta$,
$\|G_{y}(x)\|_{2}\leq\|T^{-1}\|\,\|y\|_{2}+\|T^{-1}\|\,\|F(x)-Tx\|_{2}\leq M_{0}\delta+M_{0}\cdot\frac{B}{2}r\|x\|_{2}\leq\frac{r}{2}+\frac{1}{4}\|x\|_{2}\leq r,$
so $G_{y}$ maps $B(0,r)$ to itself. Banach’s fixed-point theorem gives a unique fixed point $x^{\star}\in B(0,r)$, i.e. $F(x^{\star})=y$. The Lipschitz bound for $F^{-1}$ follows from the standard contraction argument:
$\|x^{\star}-\tilde{x}^{\star}\|_{2}=\|G_{y}(x^{\star})-G_{\tilde{y}}(\tilde{x}^{\star})\|_{2}\leq\frac{1}{2}\|x^{\star}-\tilde{x}^{\star}\|_{2}+M_{0}\|y-\tilde{y}\|_{2}.$
Rearranging yields $\|x^{\star}-\tilde{x}^{\star}\|_{2}\leq 2M_{0}\|y-\tilde{y}\|_{2}$. ∎
7.5. Quantitative inverse chart (shifted-DFT square design)
Assume $M=N$ and choose the square shifted-DFT design (Theorem˜6.13 with $M=N$):
$x_{k}:=\frac{2\pi(k-\tfrac{1}{2})}{\Lambda},\quad k=1,\dots,N,\qquad z_{k}=x_{k}+i\eta.$
Lemma 7.7 (Realification does not change smin).
Let $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ be complex-linear and identify $\mathbb{C}^{N}\simeq\mathbb{R}^{2N}$ via real/imag parts. Let $T_{\mathbb{R}}:\mathbb{R}^{2N}\to\mathbb{R}^{2N}$ be the induced real-linear map. Then
$\sigma_{\min}(T_{\mathbb{R}})=\sigma_{\min}(T).$
Proof.
One checks that $T_{\mathbb{R}}^{\top}T_{\mathbb{R}}$ is orthogonally similar to $\operatorname{diag}(T^{*}T,T^{*}T)$. Thus the eigenvalues (hence singular values) of $T_{\mathbb{R}}$ are those of $T$ repeated twice. ∎
Theorem 7.8 (Quantitative nonlinear inverse chart with explicit constants).
Assume $M=N$ and take the shifted-DFT design above. Let $T:=D\mathcal{S}(0)$ (complex-linear on $\mathbb{C}^{N}$). Let $Z:=\max_{k}|z_{k}|$ and choose $r_{d}$ by Lemma˜7.2. Let $B$ be the explicit uniform bound from Theorem˜7.5 on $B(0,r_{d})$.
Define
$\alpha$ $=\sigma_{\min}(T),$ $M_{0}$ $=\|T^{-1}\|_{2\to 2}=\frac{1}{\alpha},$
$r$ $=\min\left\{r_{d},\ \frac{1}{2BM_{0}},\ \frac{1}{2}-\varepsilon\right\},$ $\delta$ $=\frac{r}{2M_{0}}.$
Then:
  1. (i)
    (Base invertibility) One has the explicit lower bound
    $\alpha\ \geq\ 2\sqrt{N}\,\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2N}\Bigr)}\,e^{-\eta(\Lambda-\ell/2)}.$
  2. (ii)
    (Local bijectivity and stability) $\mathcal{S}$ is injective on $B(0,r)$ and for every $y$ with $\|y-\mathcal{S}(0)\|_{2}\leq\delta$ there exists a unique $\theta\in B(0,r)$ with $\mathcal{S}(\theta)=y$. Moreover the local inverse satisfies
    $\|\mathcal{S}^{-1}(y)-\mathcal{S}^{-1}(\tilde{y})\|_{2}\leq 2M_{0}\,\|y-\tilde{y}\|_{2}\quad\text{for }\|y-\mathcal{S}(0)\|_{2},\ \|\tilde{y}-\mathcal{S}(0)\|_{2}\leq\delta.$
Proof.
Item (i) is exactly the shifted-DFT lower bound in Theorem˜1.3(ii) specialized to $M=N$, since $T=D\mathcal{S}(0)$ is the Jacobian at the free point for this design.
For (ii), identify $\mathbb{C}^{N}\simeq\mathbb{R}^{2N}$ by real and imaginary parts and define the realified map
$F_{\mathbb{R}}(\theta):=\mathcal{S}_{\mathbb{R}}(\theta)-\mathcal{S}_{\mathbb{R}}(0),\qquad\theta\in B(0,r)\subset\mathbb{R}^{2N},$
where $\mathcal{S}_{\mathbb{R}}$ is the same coordinate identification applied to the output $\mathbb{C}^{N}$. Then $F_{\mathbb{R}}(0)=0$ and $DF_{\mathbb{R}}(0)=T_{\mathbb{R}}$. By Lemma˜7.7, $\sigma_{\min}(T_{\mathbb{R}})=\sigma_{\min}(T)=\alpha$, hence $\|T_{\mathbb{R}}^{-1}\|_{2\to 2}=M_{0}$.
By Theorem˜7.5, on $B(0,r_{d})$ the derivative of $\mathcal{S}$ is Lipschitz with constant $B$ (in operator norm), and therefore the same bound holds for $DF_{\mathbb{R}}$ on $B(0,r)\subset B(0,r_{d})$. By the definition of $r$ we have $BM_{0}r\leq\frac{1}{2}$. Applying Lemma˜7.6 to $F_{\mathbb{R}}$ yields: $F_{\mathbb{R}}$ is injective on $B(0,r)$, and for every $y$ with $\|y-\mathcal{S}(0)\|_{2}\leq\delta$ there exists a unique $\theta\in B(0,r)$ with $\mathcal{S}(\theta)=y$. The Lipschitz bound for the inverse map is exactly the stated estimate, translated back from the realified coordinates. ∎
7.6. Achievability: explicit reconstruction
Theorem 7.9 (Fixed-point reconstruction and stability).
Assume $M=N$ and the shifted-DFT design. Use the constants from Theorem˜7.8. Let $y$ satisfy $\|y-\mathcal{S}(0)\|_{2}\leq\delta$. Define
$\mathcal{G}_{y}(\theta):=\theta-T^{-1}(\mathcal{S}(\theta)-y),$ (66)
and iterate $\theta^{(n+1)}=\mathcal{G}_{y}(\theta^{(n)})$ starting from $\theta^{(0)}=0$. Then $\mathcal{G}_{y}$ is a contraction on $B(0,r)$, the iterates stay in $B(0,r)$, and $\theta^{(n)}\to\theta^{\star}$, where $\theta^{\star}$ is the unique solution of $\mathcal{S}(\theta)=y$ in $B(0,r)$. Moreover, if $y=\mathcal{S}(\theta_{\rm true})+e$ with $\theta_{\rm true}\in B(0,r)$ and $\|e\|_{2}\leq\delta$, then
$\|\theta^{\star}-\theta_{\rm true}\|_{2}\ \leq\ 2M_{0}\,\|e\|_{2}\ =\ \frac{2}{\alpha}\,\|e\|_{2}.$
Proof.
Differentiate Equation˜66 to obtain $D\mathcal{G}_{y}(\theta)=I-T^{-1}D\mathcal{S}(\theta)$. On $B(0,r)$, Theorem˜7.8(ii) gives $\|I-T^{-1}D\mathcal{S}(\theta)\|_{2\to 2}\leq 1/2$, hence $\mathcal{G}_{y}$ is a contraction. Convergence and stability follow from the Banach fixed-point theorem. ∎
Theorem 7.10 (Matched local stability/instability exponents near H0).
Assume the square block model $M=N$ with positivity margin $\varepsilon>0$ and fixed height $\eta>0$. Let $\mathcal{S}$ be the seam map and $T=D\mathcal{S}(0)$. Set $\alpha:=\sigma_{\min}(T)>0$ and let $B$ be a uniform bound on $\|D^{2}\mathcal{S}(\theta)\|_{2\to 2\to 2}$ on the base ball $B(0,r_{d})$.
  1. (i)
    (Achievability) On the explicit inversion radius from Theorem˜7.8, the fixed-point reconstruction from Theorem˜7.9 satisfies
    $\|\theta^{\star}-\theta_{\rm true}\|_{2}\leq\frac{2}{\alpha}\,\|e\|_{2}.$
  2. (ii)
    (Minimax lower bound, small-noise regime) If $\delta\leq\alpha^{2}/B$, then under adversarial noise $\|e\|_{2}\leq\delta$, every estimator $\widehat{\theta}$ satisfies
    $\sup_{\|e\|_{2}\leq\delta}\ \sup_{\theta\in B(0,r_{d})}\ \|\widehat{\theta}-\theta\|_{2}\ \geq\ \frac{\delta}{4\alpha}.$
  3. (iii)
    (Exponential scaling of $\alpha^{-1}$) In the block model, one always has the universal depth-barrier upper bound
    $\alpha\ \leq\ 2\sqrt{N}\,\cosh\!\Bigl(\frac{\eta\ell}{2}\Bigr)\,e^{-\eta(\Lambda-\ell/2)}.$
    Moreover, for the shifted-DFT (half-shift) square design one has the lower bound
    $\alpha\ \geq\ 2\sqrt{N}\,\sqrt{\sinh^{2}\!\Bigl(\frac{\eta\ell}{2}\Bigr)+\sin^{2}\!\Bigl(\frac{\pi}{2N}\Bigr)}\,e^{-\eta(\Lambda-\ell/2)}.$
    Consequently, under such near-optimal designs,
    $\alpha^{-1}\asymp\frac{1}{\sqrt{N}}\,e^{\eta(\Lambda-\ell/2)}$
    up to explicit multiplicative constants depending only on $\eta\ell$ and $N$ through the row-factor terms.
Proof.
Items (i) and (ii) are precisely Theorem˜7.9 and the small-noise regime in Theorem˜8.3. Item (iii) combines Theorem˜6.7 (with $M=N$) and Theorem˜6.13 (with $M=N$). ∎
8. Deterministic minimax barriers under fixed-height sampling
8.1. Two-point lower bound
Lemma 8.1 (Two-point lower bound under adversarial noise).
Let $F:\mathbb{R}^{d}\to\mathbb{R}^{m}$ and fix $\delta>0$. Consider observations $y=F(\theta)+e$ with $\|e\|_{2}\leq\delta$. If there exist $\theta_{0},\theta_{1}$ with $\|F(\theta_{1})-F(\theta_{0})\|_{2}\leq\delta$, then for any estimator $\widehat{\theta}=\widehat{\theta}(y)$,
$\sup_{\|e\|_{2}\leq\delta}\ \max\bigl\{\|\widehat{\theta}(y)-\theta_{0}\|_{2},\ \|\widehat{\theta}(y)-\theta_{1}\|_{2}\bigr\}\ \geq\ \frac{1}{2}\|\theta_{1}-\theta_{0}\|_{2}.$
Proof.
Deterministic two-point (Le Cam type) argument; see [22]. ∎
8.2. Local minimax barrier near the free point
Lemma 8.2 (Second-order remainder bound).
Assume $M=N$ and use the setup of Theorem˜7.5. On $B(0,r_{d})$, for all $\theta$,
$\|\mathcal{S}(\theta)-\mathcal{S}(0)-T\theta\|_{2}\leq\frac{B}{2}\,\|\theta\|_{2}^{2},\quad\text{where }T=D\mathcal{S}(0).$
Proof.
Taylor’s theorem with integral remainder in Banach spaces gives $\mathcal{S}(\theta)=\mathcal{S}(0)+T\theta+\int_{0}^{1}(1-t)\,D^{2}\mathcal{S}(t\theta)[\theta,\theta]\,dt$. Taking norms and using $\sup\|D^{2}\mathcal{S}\|\leq B$ on $B(0,r_{d})$ yields the claim. ∎
Theorem 8.3 (Local minimax barrier near the free point).
Assume $M=N$ and fix height $\eta>0$ and nodes $z_{k}=x_{k}+i\eta$. Let $T=D\mathcal{S}(0)$ and let $\alpha:=\sigma_{\min}(T)>0$. Assume $r_{d}$ and $B$ are as in Theorem˜7.5. Fix any noise level $\delta>0$ and set
$t:=\min\left\{r_{d},\ \frac{\delta}{2\alpha},\ \sqrt{\frac{\delta}{B}}\right\}.$
Consider the adversarial-noise observation model
$y=\mathcal{S}(\theta)+e,\qquad\|e\|_{2}\leq\delta,\qquad\theta\in B(0,r_{d}).$
Then for any estimator $\widehat{\theta}=\widehat{\theta}(y)$,
$\sup_{\|e\|_{2}\leq\delta}\ \sup_{\theta\in B(0,r_{d})}\ \|\widehat{\theta}(y)-\theta\|_{2}\ \geq\ \frac{t}{2}.$
In the regime $\delta\leq\alpha^{2}/B$,
$\sup_{\|e\|_{2}\leq\delta}\ \sup_{\theta\in B(0,r_{d})}\ \|\widehat{\theta}(y)-\theta\|_{2}\ \geq\ \frac{\delta}{4\alpha}.$
Combining with Theorem˜6.7 (with $M=N$) yields the explicit exponential lower bound
$\sup_{\|e\|_{2}\leq\delta}\ \sup_{\theta\in B(0,r_{d})}\ \|\widehat{\theta}(y)-\theta\|_{2}\ \geq\ \frac{\delta}{8\sqrt{N}\,\cosh(\eta\ell/2)}\,e^{\eta(\Lambda-\ell/2)}.$
Proof.
Let $h$ be a unit right-singular vector with $\|Th\|_{2}=\alpha$. Set $\theta_{0}=0$ and $\theta_{1}=th$. Then $\theta_{0},\theta_{1}\in B(0,r_{d})$. Using Lemma˜8.2,
$\|\mathcal{S}(\theta_{1})-\mathcal{S}(\theta_{0})\|_{2}\leq\|T\theta_{1}\|_{2}+\frac{B}{2}\|\theta_{1}\|_{2}^{2}=t\alpha+\frac{B}{2}t^{2}.$
By the choice of $t\leq\delta/(2\alpha)$ and $t\leq\sqrt{\delta/B}$, the right-hand side is $\leq\delta$. Apply Lemma˜8.1. The final exponential form uses Theorem˜6.7 to upper bound $\alpha$. ∎
9. A functional-analytic model: de Branges/Paley–Wiener identification and singular values
This section records two structural facts that explain why the block model is naturally Fourier-like. First, at $H_{0}$ the linearized map is a weighted Fourier–Laplace transform whose image is a de Branges space of Paley–Wiener type. Second, the continuum analogue of sampling stability can be phrased in terms of singular values of a time–frequency limiting operator, classically studied via prolate spheroidal wave functions.
9.1. Weighted Fourier–Laplace transform and a Paley–Wiener type space
Fix $\eta>0$ and define the weighted Fourier–Laplace transform
$(\mathcal{L}_{\eta,\Lambda}q)(x):=\int_{0}^{\Lambda}q(s)e^{-\eta s}e^{ixs}\,ds,\qquad x\in\mathbb{R}.$ (67)
Then Theorem˜3.5 can be rewritten as
$Dv_{H_{0},\Lambda}(x+i\eta)[q]=-(x+i\eta)i\,(\mathcal{L}_{\eta,\Lambda}q)(x).$ (68)
In $L^{2}$ scale, $\mathcal{L}_{\eta,\Lambda}$ is the Fourier transform of the compactly supported function $s\mapsto q(s)e^{-\eta s}\mathbf{1}_{[0,\Lambda]}(s)$ and hence produces an entire function of exponential type $\leq\Lambda$.
Proposition 9.1 (Paley–Wiener structure and de Branges identification at the free point).
Let $f\in L^{2}(0,\Lambda)$ and define
$F(z)=\int_{0}^{\Lambda}f(s)e^{izs}\,ds.$
Then $F$ is entire of exponential type $\leq\Lambda$ and $F|_{\mathbb{R}}\in L^{2}(\mathbb{R})$ (a Paley–Wiener theorem; see, e.g., [12]), and moreover there exists $G\in PW_{\Lambda/2}$ such that
$F(z)=e^{iz\Lambda/2}G(z).$
Equivalently, the range corresponds to $e^{iz\Lambda/2}PW_{\Lambda/2}$. In de Branges notation, the standard centered Paley–Wiener space is identified as
$PW_{\Lambda/2}=\mathcal{H}(E_{0}),\qquad E_{0}(z):=e^{-i\Lambda z/2};$
See [6, 13].
Proof.
Define $g\in L^{2}(\mathbb{R})$ by $g(s):=f(s)\mathbf{1}_{[0,\Lambda]}(s)$. Then
$F(z)=\int_{0}^{\Lambda}f(s)e^{izs}\,ds=\int_{\mathbb{R}}g(s)e^{izs}\,ds,\qquad z\in\mathbb{C}.$
In particular $F$ is entire and satisfies the standard Paley–Wiener growth bound (see, e.g., [11]) $|F(x+iy)|\leq e^{|y|\Lambda}\|g\|_{L^{1}(\mathbb{R})}\leq\sqrt{\Lambda}\,e^{|y|\Lambda}\|f\|_{L^{2}(0,\Lambda)}$ by Cauchy–Schwarz. Moreover, by the Paley–Wiener theorem under the Fourier normalization used in this paper (equivalently, by Plancherel together with $\operatorname{supp}g\subset[0,\Lambda]$; see, e.g., [12]), one has $F|_{\mathbb{R}}\in L^{2}(\mathbb{R})$ and $F$ has exponential type $\leq\Lambda$.
For the centered identification, set $\widetilde{g}(t):=g(t+\Lambda/2)=f(t+\Lambda/2)\mathbf{1}_{[-\Lambda/2,\Lambda/2]}(t)$ and define
$G(z):=\int_{-\Lambda/2}^{\Lambda/2}\widetilde{g}(t)e^{izt}\,dt.$
Then $G\in PW_{\Lambda/2}$ and $F(z)=e^{iz\Lambda/2}G(z)$ by the change of variables $s=t+\Lambda/2$. ∎
9.2. Sampling as a kernel analysis operator
Let $PW_{\Lambda/2}$ be equipped with its standard Hilbert structure. Evaluation at a point $w\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ is continuous and represented by the reproducing kernel $k_{w}$. Sampling on a horizontal line corresponds to applying the analysis operator of a family of kernels.
Definition 9.2 (Kernel sampling operator on a fixed height).
Fix $\eta>0$ and nodes $x_{1},\dots,x_{M}\in\mathbb{R}$. Define $\mathrm{Ev}_{\eta,x}$ on $PW_{\Lambda/2}$ by
$\mathrm{Ev}_{\eta,x}(F):=(F(x_{k}+i\eta))_{k=1}^{M}\in\mathbb{C}^{M}.$
Proposition 9.3 (Frame bounds and singular values).
Let $\mathcal{H}$ be a reproducing kernel Hilbert space and let $w_{1},\dots,w_{M}$ be points in its domain with kernels $k_{w_{k}}$. Let $A$ be the $M\times M$ Gram matrix $A_{jk}:=\langle k_{w_{k}},k_{w_{j}}\rangle_{\mathcal{H}}$. Then the nonzero singular values of $\mathrm{Ev}:F\mapsto(F(w_{k}))_{k=1}^{M}$ restricted to $\mathrm{span}\{k_{w_{k}}\}$ are the square roots of the nonzero eigenvalues of $A$. In particular, kernel frame lower bounds are equivalent to quantitative lower bounds on the smallest singular value.
Proof.
Throughout, $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ is linear in the first argument and conjugate–linear in the second, and $\langle a,b\rangle_{\mathbb{C}^{M}}:=\sum_{j=1}^{M}a_{j}\overline{b_{j}}$. Let $\mathrm{Ev}_{w}:\mathcal{H}\to\mathbb{C}^{M}$ be $\mathrm{Ev}_{w}f=(f(w_{j}))_{j=1}^{M}$. Since $\mathcal{H}$ is an RKHS, $f(w_{j})=\langle f,k_{w_{j}}\rangle_{\mathcal{H}}$, hence for $c\in\mathbb{C}^{M}$,
$\langle\mathrm{Ev}_{w}f,c\rangle_{\mathbb{C}^{M}}=\sum_{j=1}^{M}f(w_{j})\overline{c_{j}}=\sum_{j=1}^{M}\langle f,k_{w_{j}}\rangle_{\mathcal{H}}\overline{c_{j}}=\Big\langle f,\sum_{j=1}^{M}c_{j}k_{w_{j}}\Big\rangle_{\mathcal{H}}.$
Therefore $\mathrm{Ev}_{w}^{*}c=\sum_{j=1}^{M}c_{j}k_{w_{j}}$.
Now compute $\mathrm{Ev}_{w}\mathrm{Ev}_{w}^{*}:\mathbb{C}^{M}\to\mathbb{C}^{M}$: for $c\in\mathbb{C}^{M}$ and each $j$,
$(\mathrm{Ev}_{w}\mathrm{Ev}_{w}^{*}c)_{j}=(\mathrm{Ev}_{w}(\sum_{k=1}^{M}c_{k}k_{w_{k}}))_{j}=\sum_{k=1}^{M}c_{k}\,k_{w_{k}}(w_{j})=\sum_{k=1}^{M}A_{jk}c_{k},$
where $A_{jk}:=k_{w_{k}}(w_{j})=\langle k_{w_{k}},k_{w_{j}}\rangle_{\mathcal{H}}$ is the Gram matrix. Hence $\mathrm{Ev}_{w}\mathrm{Ev}_{w}^{*}=A$ as an operator on $\mathbb{C}^{M}$; in particular $A$ is Hermitian positive semidefinite.
By the singular value identity, the nonzero eigenvalues of $\mathrm{Ev}_{w}\mathrm{Ev}_{w}^{*}$ are the squares of the nonzero singular values of $\mathrm{Ev}_{w}$, with matching multiplicities. Since $\mathrm{Ev}_{w}\mathrm{Ev}_{w}^{*}=A$, the claim follows. ∎
9.3. Continuum singular values via time–frequency limiting
A continuum analogue of sampling stability is governed by singular values of a time–frequency limiting operator. Let $\Omega>0$ and define the (unweighted) truncated Fourier transform
$(\mathcal{F}_{\Lambda,\Omega}f)(\xi):=\int_{0}^{\Lambda}f(s)e^{-i\xi s}\,ds,\qquad\xi\in[-\Omega,\Omega].$
Then $\mathcal{F}_{\Lambda,\Omega}:L^{2}(0,\Lambda)\to L^{2}([-\Omega,\Omega])$ is compact and $\mathcal{F}_{\Lambda,\Omega}^{*}\mathcal{F}_{\Lambda,\Omega}$ is the classical time–frequency limiting operator whose eigenfunctions are prolate spheroidal wave functions; see [20, 15, 16].
Proposition 9.4 (Classical singular value picture).
Let $\sigma_{1}\geq\sigma_{2}\geq\cdots$ be the singular values of $\mathcal{F}_{\Lambda,\Omega}$. There is an effective dimension (the Shannon number) $N_{0}\asymp\Lambda\Omega/\pi$ such that $\sigma_{n}\approx 1$ for $n\lesssim N_{0}$, while $\sigma_{n}$ decays rapidly for $n\gtrsim N_{0}$, with a sharp transition described by the asymptotic theory of [16] (see also [20]).
Proof.
This is a classical theorem in time–frequency analysis; the operator $\mathcal{F}_{\Lambda,\Omega}^{*}\mathcal{F}_{\Lambda,\Omega}$ is an integral operator with the sinc kernel and its eigenstructure is the prolate spheroidal theory developed in [15, 20, 16]. ∎
The weighted transform in Equation˜67 is obtained by composing $\mathcal{F}_{\Lambda,\Omega}$ with multiplication by $e^{-\eta s}$ on $0,\Lambda]$. Since $0\leq e^{-\eta s}\leq 1$, the singular values of the weighted operator are bounded above by those of the unweighted one. This provides a continuum analogue of the depth barrier, now as an intrinsic compactness mechanism: fixed-height data corresponds to sampling entire functions away from the boundary, hence it necessarily suppresses high time–frequency content.
10. A base-point-free exponential barrier from spectral smoothing
This section is independent of any base point. It shows that fixed-height seam data necessarily encodes a Poisson-smoothed version of the spectral measure, and therefore any attempt to recover unsmoothed spectral features from finitely many samples is exponentially ill-conditioned in the feature frequency.
10.1. de Branges correspondence and the effect of fixing a free tail
Theorem 10.1 (de Branges correspondence (trace-normed surjectivity onto Herglotz functions)).
Let $m$ be a Herglotz–Nevanlinna function on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (analytic and $\operatorname{Im}m(z)\geq 0$ for $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$). Then there exists a trace-normed Hamiltonian $H$ on $0,\infty)$ such that $m$ is the Weyl coefficient of the associated canonical system. In the trace-normed gauge, $H$ is unique up to equality a.e.
Proof.
This is the de Branges correspondence in the trace-normed class: the map $H\mapsto m_{H}$ is a bijection between trace-normed half-line Hamiltonians and Herglotz–Nevanlinna functions, with uniqueness up to equality a.e.; see [9, Thm. 2.2]. ∎
Theorem 10.2 (Free-tail Weyl functions are meromorphic of controlled exponential type).
Let $H$ be trace-normed on $0,\Lambda]$ and extended by the free tail $H(s)=\tfrac{1}{2}I$ for $s\geq\Lambda$. Write
$\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}A_{H}(z)&B_{H}(z)\\ C_{H}(z)&D_{H}(z)\end{pmatrix}$.
Then:
  1. (i)
    each entry $A_{H},B_{H},C_{H},D_{H}$ is entire and satisfies the global bound
    $\|\Phi_{H}(\Lambda,0;z)\|_{\mathrm{op}}\leq\exp(\Lambda|z|),$
    hence is of exponential type at most $\Lambda$;
  2. (ii)
    the Weyl function $m_{H,\Lambda}$ and the Schur function $v$ extend meromorphically to $\mathbb{C}$ via $m(z)=(iA_{H}(z)-C_{H}(z))/(D_{H}(z)-iB_{H}(z))$ and v(z)=(m(z)i)/(m(z)+i)v(z)=(m(z)-i)/(m(z)+i);
  3. (iii)
    $m$ has no poles in $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (only removable singularities there), so all poles of the meromorphic continuation lie in $\{z:\operatorname{Im}z<0\}\cup\mathbb{R}$.
Proof.
By the Volterra equation Equation˜22 and Picard iteration, $z\mapsto\Phi_{H}(\Lambda,0;z)$ is entire. Moreover, since $H(s)\succeq 0$ and $\operatorname{tr}H(s)=1$, one has $\|H(s)\|_{\mathrm{op}}\leq 1$ a.e., and $\|\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\|_{\mathrm{op}}=1$. Taking operator norms in Equation˜22 yields
$\|\Phi_{H}(s,0;z)\|_{\mathrm{op}}\leq 1+|z|\int_{0}^{s}\|\Phi_{H}(u,0;z)\|_{\mathrm{op}}\,du,$
so Grönwall gives $\|\Phi_{H}(\Lambda,0;z)\|_{\mathrm{op}}\leq e^{|z|\Lambda}$ for all $z\in\mathbb{C}$, proving (i). Statement (ii) is immediate from the transfer-matrix Weyl formula Equation˜6 and the fact that $A_{H},B_{H},C_{H},D_{H}$ are entire and $D_{H}-iB_{H}\not\equiv 0$. Finally, (iii) follows from Equation˜7. ∎
Remark 10.3 (A fixed free tail is a strict subclass).
If one additionally imposes the free tail $H(s)\equiv\tfrac{1}{2}I$ for all $s\geq\Lambda$, then the class of realizable Weyl functions is strictly smaller than the class of all Herglotz functions.
Indeed, the free tail pins down the Weyl datum at $s=\Lambda$ to $m_{\Lambda}(z)\equiv i$ (equivalently $v_{\Lambda}(z)=0$). For a canonical system on the finite interval $0,\Lambda]$ with fundamental matrix $\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}a(z)&b(z)\\ c(z)&d(z)\end{pmatrix}$, the Weyl function at the left endpoint is obtained from the right-end Weyl datum by a Möbius transformation:
$m_{0}(z)=\frac{a(z)\,m_{\Lambda}(z)+b(z)}{c(z)\,m_{\Lambda}(z)+d(z)}.$ (69)
This follows directly from the transfer relation $Y(\Lambda;z)=\Phi_{H}(\Lambda,0;z)Y(0;z)$ and the definition of endpoint Weyl data. Since $\Phi_{H}(\Lambda,0;z)$ depends analytically on $z$ for finite $\Lambda$, (69) shows $m_{0}$ is a quotient of entire functions, hence meromorphic on $\mathbb{C}$.
But not every Herglotz function is meromorphic on $\mathbb{C}$. For an explicit counterexample, take
$m(z):=\int_{-1}^{1}\frac{dt}{t-z}=\log\!\Bigl(\frac{1-z}{-1-z}\Bigr),$
which is Herglotz on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ but has logarithmic branch points at $\pm 1$ (for the principal branch with a cut along [1,1][-1,1]), hence is not meromorphic on $\mathbb{C}$. Therefore imposing a free tail after $\Lambda$ cuts down the realizable class strictly.
10.2. A transfer-matrix characterization of free-tail Weyl functions
The free-tail condition $H(s)\equiv\tfrac{1}{2}I$ for $s\geq\Lambda$ singles out a proper subclass of Herglotz functions, but it also admits a clean intrinsic description in terms of entire transfer matrices.
Definition 10.4 (J-contractive entire matrices).
Let $\Phi:\mathbb{C}\to\mathbb{C}^{2\times 2}$ be entire with $\det\Phi\equiv 1$. We say that $\Phi$ is $J$-contractive in $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (for our sign convention $JY^{\prime}=zHY$) if
$i\big(\Phi(z)^{*}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Phi(z)-\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\big)\ \succeq\ 0,\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\},$ (70)
and $J$-unitary on $\mathbb{R}$ if $\Phi(x)^{\top}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,\Phi(x)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$ for a.e. $x\in\mathbb{R}$.
Theorem 10.5 (Inverse monodromy for trace-normed canonical systems).
Fix $\Lambda>0$. Let $\Phi:\mathbb{C}\to\mathbb{C}^{2\times 2}$ be an entire matrix function with $\Phi(0)=I$ and $\det\Phi\equiv 1$. Assume:
  1. (i)
    $\Phi$ is $J$-contractive in $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and $J$-unitary on $\mathbb{R}$ in the sense of Definition 10.4;
  2. (ii)
    (finite growth) there is a constant $C\geq 1$ such that $\|\Phi(z)\|_{\mathrm{op}}\leq Ce^{\Lambda|z|}$ for all $z\in\mathbb{C}$.
Then there exists a trace-normed Hamiltonian $H$ on $0,\Lambda]$, unique a.e., such that
$\Phi(z)=\Phi_{H}(\Lambda,0;z)\qquad\text{for all }z\in\mathbb{C}.$
Proof.
This is a standard realization (inverse transfer-matrix/monodromy) theorem in the Potapov–de Branges theory of canonical systems. One may reduce to the sign convention of [18, 6] by replacing $z$ with z-z, and then apply the characterization of transfer matrices as entire $J$-inner (equivalently, $J$-contractive in $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and $J$-unitary on $\mathbb{R}$) functions of finite exponential type; see, for example, [18]. ∎
Theorem 10.6 (Characterization of free-tail Weyl functions).
Fix $\Lambda>0$. Let $m$ be a Herglotz–Nevanlinna function on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. The following are equivalent:
  1. (a)
    There exists a trace-normed Hamiltonian $H$ on $0,\Lambda]$ such that the tail-extended Hamiltonian $\widehat{H}(s):=H(s)\mathbf{1}_{[0,\Lambda]}(s)+\tfrac{1}{2}I\,\mathbf{1}_{(\Lambda,\infty)}(s)$ has Weyl function $m_{\widehat{H}}(z)=m(z)$.
  2. (b)
    There exists an entire matrix function
    $\Phi(z)=\begin{pmatrix}A(z)&B(z)\\ C(z)&D(z)\end{pmatrix}$
    satisfying the hypotheses of Theorem 10.5 and such that $D(z)-iB(z)\neq 0$ for all $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, with
    $m(z)=\frac{iA(z)-C(z)}{D(z)-iB(z)},\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}.$ (71)
In either case, $H$ is determined uniquely a.e. on $0,\Lambda]$.
Proof.
[…]
Uniqueness of $H$ in (a) follows from the uniqueness statement in Theorem 10.5. ∎
Remark 10.7 (Nonuniqueness of the cutoff parameter).
If $H$ is trace-normed on $0,\Lambda]$ and free for $s\geq\Lambda$, then it is also free for $s\geq\Lambda+L$ for any $L>0$, hence the same Weyl function admits realizations with arbitrarily large cutoffs. Explicitly, appending a free segment of length $L$ multiplies both the numerator and denominator in (71) by $e^{izL/2}$, leaving $m$ unchanged (cf. the eigenvector computation in the proof of Theorem˜2.3). Thus the minimal cutoff is the meaningful invariant; it can be expressed in terms of the (mean) exponential type of the corresponding de Branges data, see [18].
10.3. Herglotz representation and Poisson smoothing
Theorem 10.8 (Herglotz representation and spectral measure).
For each fixed (tail-extended) Hamiltonian, the Weyl function $m_{H,\Lambda}$ is Herglotz–Nevanlinna and admits
$m_{H,\Lambda}(z)=c_{1}+c_{2}z+\int_{\mathbb{R}}\Bigl(\frac{1}{t-z}-\frac{t}{1+t^{2}}\Bigr)\,d\rho(t),\qquad z\in\mathbb{C}\setminus\mathbb{R},$ (72)
where $c_{1}\in\mathbb{R}$, $c_{2}\geq 0$, and $\rho$ is nondecreasing with $\int_{\mathbb{R}}\frac{d\rho(t)}{1+t^{2}}<\infty$.
Proof.
This is standard for canonical systems; see, e.g., [9]. ∎
Corollary 10.9 (Poisson smoothing).
Let $\eta>0$ and $x\in\mathbb{R}$. Writing $z=x+i\eta$, the imaginary part satisfies
$\operatorname{Im}\,m_{H,\Lambda}(x+i\eta)=c_{2}\,\eta+\int_{\mathbb{R}}\frac{\eta}{(t-x)^{2}+\eta^{2}}\,d\rho(t),$
i.e. $\operatorname{Im}m(\cdot+i\eta)$ is the Poisson integral (up to $c_{2}\eta$) of the spectral measure. Equivalently, with the normalized Poisson kernel $P_{\eta}(u):=\frac{1}{\pi}\frac{\eta}{u^{2}+\eta^{2}}$, one has
$\frac{1}{\pi}\Big(\operatorname{Im}m_{H,\Lambda}(x+i\eta)-c_{2}\eta\Big)=\int_{\mathbb{R}}P_{\eta}(x-t)\,d\rho(t).$
In particular, if $c_{2}=0$ and $d\rho(t)=f(t)\,dt$, then $\frac{1}{\pi}\operatorname{Im}m_{H,\Lambda}(x+i\eta)=(\mathcal{P}_{\eta}f)(x)$.
Proof.
Take imaginary parts in Equation˜72 and use $\operatorname{Im}\frac{1}{t-(x+i\eta)}=\frac{\eta}{(t-x)^{2}+\eta^{2}}$. ∎
Corollary 10.10 (Realizability of smooth adversarial spectral densities).
Fix $\chi\in C_{c}^{\infty}(\mathbb{R})$ and $K\geq 1$. Let $f_{0}\in C_{c}^{\infty}(\mathbb{R})$ satisfy $f_{0}\geq 0$ on $\mathbb{R}$, and choose $\varepsilon>0$ so that
$f_{0}(t)\ \geq\ \varepsilon\|\chi\|_{L^{\infty}(\mathbb{R})}\qquad\text{for all }t\in\operatorname{supp}\chi.$
Define
$f^{\pm}(t):=f_{0}(t)\pm\varepsilon\cos(Kt)\chi(t).$
Then $f^{\pm}\geq 0$ and $f^{\pm}\in L^{1}(\mathbb{R})$. Moreover the functions
$m_{\pm}(z):=\int_{\mathbb{R}}\Bigl(\frac{1}{t-z}-\frac{t}{1+t^{2}}\Bigr)f^{\pm}(t)\,dt,\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\},$
are Herglotz–Nevanlinna. Hence each $m_{\pm}$ is the Weyl coefficient of a (unique trace-normed) canonical system on $0,\infty)$. Moreover, the boundary densities satisfy $f^{\pm}(x)=\frac{1}{\pi}\operatorname{Im}m_{\pm}(x+i0)$ for a.e. $x\in\mathbb{R}$.
Proof.
Since $f_{0},\chi\in C_{c}^{\infty}(\mathbb{R})$ we have $f^{\pm}\in C_{c}^{\infty}(\mathbb{R})\subset L^{1}(\mathbb{R})$, and also $\int\frac{f^{\pm}(t)}{1+t^{2}}\,dt<\infty$. Thus the integral defining $m_{\pm}(z)$ converges absolutely for each $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$.
For nonnegativity: for $t\in\operatorname{supp}\chi$,
$|\,\varepsilon\cos(Kt)\chi(t)\,|\leq\varepsilon|\chi(t)|\leq\varepsilon\|\chi\|_{\infty}\leq f_{0}(t),$
hence $f^{\pm}(t)\geq 0$ on $\operatorname{supp}\chi$. Outside $\operatorname{supp}\chi$ we have $f^{\pm}=f_{0}\geq 0$.
Now write $z=x+i\eta$ with $\eta>0$. The subtraction term $\int\frac{t}{1+t^{2}}f^{\pm}(t)\,dt$ is real, hence does not affect imaginary parts. Also
$\operatorname{Im}\Bigl(\frac{1}{t-(x+i\eta)}\Bigr)=\frac{\eta}{(t-x)^{2}+\eta^{2}}>0.$
Therefore
$\operatorname{Im}\,m_{\pm}(x+i\eta)=\int_{\mathbb{R}}\frac{\eta}{(t-x)^{2}+\eta^{2}}\,f^{\pm}(t)\,dt\ \geq\ 0,$
so $m_{\pm}$ are Herglotz–Nevanlinna. Finally apply Theorem˜10.1. ∎
10.4. From seam samples to Poisson-smoothed spectral data
Since $v$ is the Cayley transform of $m$, one can recover $m$ from $v$ pointwise by the inverse Cayley map
$m(z)=i\,\frac{1+v(z)}{1-v(z)}.$ (73)
In particular, $\operatorname{Im}m(z)$ can be written explicitly in terms of $v(z)$:
$\operatorname{Im}m(z)=\frac{1-|v(z)|^{2}}{|1-v(z)|^{2}},\qquad z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}.$ (74)
Thus fixed-height seam measurements $v(x_{k}+i\eta)$ determine samples of the Poisson smoothing in Corollary˜10.9.
10.5. A quantitative minimax obstruction for de-smoothing
Fourier convention. In this subsection we use the unitary Fourier transform on $\mathbb{R}$ (so that Plancherel holds without $2\pi$ factors). This choice affects only numerical constants in the Poisson-smoothing estimates.
For $\eta>0$, recall the Poisson kernel (as in Corollary˜10.9)
$P_{\eta}(x):=\frac{1}{\pi}\frac{\eta}{x^{2}+\eta^{2}},\qquad(\mathcal{P}_{\eta}f)(x):=(P_{\eta}*f)(x).$
With the unitary Fourier transform
$\widehat{f}(\xi)=(2\pi)^{-1/2}\int_{\mathbb{R}}f(t)e^{-it\xi}\,dt,\qquad\check{g}(t)=(2\pi)^{-1/2}\int_{\mathbb{R}}g(\xi)e^{it\xi}\,d\xi,$
one has the convolution identity $\widehat{f*g}=(2\pi)^{1/2}\,\widehat{f}\,\widehat{g}$ and the explicit transform
$\widehat{P_{\eta}}(\xi)=(2\pi)^{-1/2}e^{-\eta|\xi|}.$
See, e.g., [11] for these Fourier and Poisson kernel identities. Consequently,
$\widehat{\mathcal{P}_{\eta}f}(\xi)=\widehat{P_{\eta}*f}(\xi)=(2\pi)^{1/2}\widehat{P_{\eta}}(\xi)\widehat{f}(\xi)=e^{-\eta|\xi|}\widehat{f}(\xi).$
Lemma 10.11 (Exponential Linfty damping for spectrally localized high-frequency perturbations).
Fix $\eta>0$. Let $\varphi\in C_{c}^{\infty}([-1,1])$ be real and even, and let $h:=\check{\varphi}$, so $h\in\mathcal{S}(\mathbb{R})$ is real and even. For integers $K\geq 2$, define
$h_{K}(t):=h(t)\cos(Kt).$
Then $h_{K}\in\mathcal{S}(\mathbb{R})$, and
$\|\mathcal{P}_{\eta}h_{K}\|_{L^{\infty}(\mathbb{R})}\leq(2\pi)^{-1/2}\,\|\varphi\|_{L^{1}(\mathbb{R})}\,e^{-\eta(K-1)}.$ (75)
Proof.
Because $\varphi$ is real and even, $h=\check{\varphi}$ is real and even. Using $\cos(Kt)=\tfrac{1}{2}(e^{iKt}+e^{-iKt})$ and the modulation rule for the unitary Fourier transform,
$\widehat{h_{K}}(\xi)=\frac{1}{2}\,\widehat{h}(\xi-K)+\frac{1}{2}\,\widehat{h}(\xi+K)=\frac{1}{2}\,\varphi(\xi-K)+\frac{1}{2}\,\varphi(\xi+K).$
Hence $\operatorname{supp}\widehat{h_{K}}\subset[K-1,K+1]\cup[-K-1,-K+1]$, so on $\operatorname{supp}\widehat{h_{K}}$ we have $|\xi|\geq K-1$ and thus $e^{-\eta|\xi|}\leq e^{-\eta(K-1)}$. Therefore,
$\|\widehat{\mathcal{P}_{\eta}h_{K}}\|_{L^{1}}=\int_{\mathbb{R}}e^{-\eta|\xi|}\,|\widehat{h_{K}}(\xi)|\,d\xi\leq e^{-\eta(K-1)}\|\widehat{h_{K}}\|_{L^{1}}.$
Moreover,
$\|\widehat{h_{K}}\|_{L^{1}}\leq\tfrac{1}{2}\|\varphi(\cdot-K)\|_{L^{1}}+\tfrac{1}{2}\|\varphi(\cdot+K)\|_{L^{1}}=\|\varphi\|_{L^{1}}.$
Finally, since $g=\check{\widehat{g}}$ and $\|\check{u}\|_{\infty}\leq(2\pi)^{-1/2}\|u\|_{1}$ for the unitary transform,
$\|\mathcal{P}_{\eta}h_{K}\|_{\infty}\leq(2\pi)^{-1/2}\,\|\widehat{\mathcal{P}_{\eta}h_{K}}\|_{1}\leq(2\pi)^{-1/2}\,\|\varphi\|_{1}\,e^{-\eta(K-1)}.$
Theorem 10.12 (De-smoothing from finitely many fixed-height samples is exponentially ill-conditioned).
Fix $\eta>0$ and sample points $x_{1},\dots,x_{M}\in\mathbb{R}$. Let $\varphi\in C_{c}^{\infty}([-1,1])$ be real and even, set $h=\check{\varphi}\in\mathcal{S}(\mathbb{R})$, and define $h_{K}(t)=h(t)\cos(Kt)$.
Let $\psi(t):=(1+t^{2})^{-2}$ and set
$R:=\sup_{t\in\mathbb{R}}\frac{|h(t)|}{\psi(t)}<\infty.$
Fix $\varepsilon>0$ and define
$f_{0}(t):=2\varepsilon R\,\psi(t),\qquad f^{\pm}(t):=f_{0}(t)\pm\varepsilon h_{K}(t),\qquad K\geq 2.$
Then $f^{\pm}\geq 0$ on $\mathbb{R}$ and $\int_{\mathbb{R}}\frac{f^{\pm}(t)}{1+t^{2}}\,dt<\infty$. Let $g_{f}:=\mathcal{P}_{\eta}f$. For the two alternatives $f^{\pm}$ one has
$\|(g_{f^{+}}(x_{k})-g_{f^{-}}(x_{k}))_{k=1}^{M}\|_{2}\leq c_{1}\,\varepsilon\,e^{-\eta(K-1)},\qquad c_{1}:=2\sqrt{M}\,(2\pi)^{-1/2}\|\varphi\|_{L^{1}}.$
Consequently, for noisy samples
$y_{k}=g_{f}(x_{k})+e_{k},\qquad\|(e_{k})_{k=1}^{M}\|_{2}\leq\delta,$
if $\delta\geq\frac{1}{2}c_{1}\varepsilon e^{-\eta(K-1)}$, then for every estimator $\widehat{f}=\widehat{f}(y)$,
$\sup_{\|e\|_{2}\leq\delta}\ \max_{\pm}\ \|\widehat{f}-f^{\pm}\|_{L^{2}(\mathbb{R})}\ \geq\ \varepsilon\,\|h_{K}\|_{L^{2}(\mathbb{R})}.$
Moreover, there exists $K_{0}\geq 2$ (depending only on $h$) such that for all $K\geq K_{0}$,
$\|h_{K}\|_{L^{2}(\mathbb{R})}\geq\frac{1}{2}\|h\|_{L^{2}(\mathbb{R})},$
and hence for all $K\geq K_{0}$,
$\sup_{\|e\|_{2}\leq\delta}\ \max_{\pm}\ \|\widehat{f}-f^{\pm}\|_{L^{2}(\mathbb{R})}\ \geq\ \frac{\varepsilon}{2}\,\|h\|_{L^{2}(\mathbb{R})}.$
Proof.
Positivity and integrability. Since $h\in\mathcal{S}(\mathbb{R})$ and $\psi(t)=(1+t^{2})^{-2}>0$, the ratio $|h(t)|/\psi(t)$ is continuous and tends to 0 as $|t|\to\infty$, hence $R<\infty$. Then for all $t$,
$|h_{K}(t)|\leq|h(t)|\leq R\psi(t),$
so
$f^{\pm}(t)=2\varepsilon R\psi(t)\pm\varepsilon h_{K}(t)\geq 2\varepsilon R\psi(t)-\varepsilon|h_{K}(t)|\geq\varepsilon R\psi(t)\geq 0.$
Also $\psi(t)/(1+t^{2})=(1+t^{2})^{-3}\in L^{1}(\mathbb{R})$ and $h_{K}\in\mathcal{S}(\mathbb{R})\subset L^{1}(\mathbb{R})$, hence $\int_{\mathbb{R}}\frac{f^{\pm}(t)}{1+t^{2}}\,dt<\infty$.
Exponential smallness of the sample difference. Linearity gives $g_{f^{+}}-g_{f^{-}}=2\varepsilon\,\mathcal{P}_{\eta}h_{K}$. Thus
$\|(g_{f^{+}}(x_{k})-g_{f^{-}}(x_{k}))_{k=1}^{M}\|_{2}\leq 2\varepsilon\sqrt{M}\,\|\mathcal{P}_{\eta}h_{K}\|_{\infty}.$
Apply Lemma 10.11 to obtain the stated bound with $c_{1}$.
Two-point lower bound under adversarial noise. Let $\Delta:=(g_{f^{+}}(x_{k})-g_{f^{-}}(x_{k}))_{k=1}^{M}$. If $\delta\geq\|\Delta\|_{2}/2$, then the single observation
$y:=(g_{f^{+}}(x_{k}))_{k=1}^{M}-\Delta/2$
is feasible under both alternatives using noises $e^{+}:=-\Delta/2$ and $e^{-}:=+\Delta/2$, each of norm $\leq\delta$. Hence for any estimator $\widehat{f}(y)$, the triangle inequality yields
$\|f^{+}-f^{-}\|_{2}\leq\|\widehat{f}-f^{+}\|_{2}+\|\widehat{f}-f^{-}\|_{2}\leq 2\max_{\pm}\|\widehat{f}-f^{\pm}\|_{2},$
so
$\max_{\pm}\|\widehat{f}-f^{\pm}\|_{L^{2}(\mathbb{R})}\geq\frac{1}{2}\|f^{+}-f^{-}\|_{L^{2}(\mathbb{R})}=\varepsilon\|h_{K}\|_{L^{2}(\mathbb{R})}.$
A uniform-in-$K$ lower bound in the high-frequency regime. Since $h\in\mathcal{S}(\mathbb{R})$, we have $h^{2}\in L^{1}(\mathbb{R})$. Using $\cos^{2}(Kt)=\tfrac{1}{2}(1+\cos(2Kt))$,
$\|h_{K}\|_{2}^{2}=\int_{\mathbb{R}}h(t)^{2}\cos^{2}(Kt)\,dt=\frac{1}{2}\|h\|_{2}^{2}+\frac{1}{2}\int_{\mathbb{R}}h(t)^{2}\cos(2Kt)\,dt.$
By the Riemann–Lebesgue lemma, $\int h(t)^{2}\cos(2Kt)\,dt\to 0$ as $K\to\infty$. Therefore there exists $K_{0}$ such that for all $K\geq K_{0}$,
$\left|\int_{\mathbb{R}}h(t)^{2}\cos(2Kt)\,dt\right|\leq\frac{1}{2}\|h\|_{2}^{2},$
and hence
$\|h_{K}\|_{2}^{2}\geq\frac{1}{2}\|h\|_{2}^{2}-\frac{1}{4}\|h\|_{2}^{2}=\frac{1}{4}\|h\|_{2}^{2},\qquad K\geq K_{0}.$
Taking square-roots yields $\|h_{K}\|_{2}\geq\frac{1}{2}\|h\|_{2}$ for all $K\geq K_{0}$. ∎
10.6. Hardy space on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ and Carleson measures
Let $H^{2}(\{z\in\mathbb{C}:\ \operatorname{Im}z>0\})$ denote the Hardy space on the upper half-plane, equipped with the norm
$\|F\|_{H^{2}(\{z\in\mathbb{C}:\ \operatorname{Im}z>0\})}^{2}:=\sup_{y>0}\ \int_{\mathbb{R}}|F(x+iy)|^{2}\,dx.$
Then $H^{2}(\{z\in\mathbb{C}:\ \operatorname{Im}z>0\})$ is a reproducing kernel Hilbert space; one normalization of the kernel is
$k_{w}(z):=\frac{1}{2\pi i}\,\frac{1}{\overline{w}-z},\qquad z,w\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\},$ (76)
so that $F(w)=\langle F,k_{w}\rangle_{H^{2}}$ and $k_{w}(w)=\frac{1}{4\pi\,\operatorname{Im}w}$; see [10], [8], and [14]. For an interval $I\subset\mathbb{R}$, write the Carleson box
$S(I):=\{x+iy:\ x\in I,\ 0<y<|I|\}.$
Theorem 10.13 (Carleson embedding theorem on C+).
Let $\mu$ be a positive Borel measure on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$. The embedding
$H^{2}(\{z\in\mathbb{C}:\ \operatorname{Im}z>0\})\hookrightarrow L^{2}(\mu),\qquad F\mapsto F|_{\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}},$
is bounded if and only if $\mu$ is a Carleson measure, i.e. there is $C_{\mu}<\infty$ such that
$\sup_{I\subset\mathbb{R}}\ \frac{\mu(S(I))}{|I|}\ \leq\ C_{\mu}.$ (77)
Moreover, the best embedding constant is comparable to $C_{\mu}^{1/2}$ (with universal constants).
Proof.
This is the classical Carleson theorem [3]; see also [10, 8]. ∎
Corollary 10.14 (Density criterion for discrete sampling on a fixed horizontal line).
Fix $\eta>0$ and a discrete set $X\subset\mathbb{R}$ and consider the discrete measure
$\mu_{X,\eta}:=\sum_{x\in X}\delta_{x+i\eta}.$
Then $\mu_{X,\eta}$ is Carleson if and only if
$\sup_{I\subset\mathbb{R}:\ |I|\geq\eta}\ \frac{\#(X\cap I)}{|I|}\ <\ \infty.$ (78)
In this case, the Carleson constant satisfies $C_{\mu_{X,\eta}}\asymp\sup_{|I|\geq\eta}\#(X\cap I)/|I|$.
Proof.
A point $x+i\eta$ lies in $S(I)$ if and only if $x\in I$ and $\eta<|I|$. Thus $\mu_{X,\eta}(S(I))=\#(X\cap I)$ for $|I|>\eta$ and is 0 otherwise. Insert into (77). ∎
10.7. Model spaces and de Branges–Rovnyak kernels
For a Schur function $v$ on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$, the associated de Branges–Rovnyak space $\mathcal{H}(v)$ is the reproducing kernel Hilbert space with kernel
$k_{w}^{v}(z)=\frac{1}{2\pi i}\,\frac{1-v(z)\overline{v(w)}}{\overline{w}-z}.$
This viewpoint is natural for seam sampling because $v$ itself is a Schur function and fixed-height seam measurements are evaluations of $v$. For background see [19, 21].
Remark 10.15 (Harmonic tight frames and DFT sampling).
The tight-frame condition $F_{x}^{*}F_{x}=MI_{N}$ in the block model is an instance of a harmonic finite unit-norm tight frame. General theory of finite normalized tight frames and frame potentials can be found in [2] (see also [4]).
11. An operator-theoretic corollary: compact perturbations of the free Dirac system
A trace-normed Hamiltonian $H$ on $0,\infty)$ defines a weighted Hilbert space $L^{2}_{H}:=L^{2}\bigl([0,\infty);\mathbb{C}^{2};H(s)\,ds\bigr)$ with inner product $\langle y_{1},y_{2}\rangle_{L^{2}_{H}}:=\int_{0}^{\infty}y_{1}(s)^{*}H(s)y_{2}(s)\,ds$. The canonical system
$\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,y^{\prime}(s)=z\,H(s)\,y(s)$
is the eigenvalue equation for the (formally) first-order differential expression
$\tau_{H}y:=-\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,y^{\prime}(s)$
acting in $L^{2}_{H}$. This is the standard Dirac-type interpretation of canonical systems; the Weyl function $m_{H,\Lambda}$ is exactly the Weyl–Titchmarsh function of $\tau_{H}$ for the boundary condition at $s=0$ and the $L^{2}_{H}$ condition at $\infty$. (For background, see de Branges theory and modern accounts such as [9, 18].)
If $H$ has a free tail (i.e. $H(s)=\tfrac{1}{2}I$ for $s\geq\Lambda$), then on $(\Lambda,\infty)$ the dynamics is that of the free Dirac system, and $\tau_{H}$ may be viewed as a compactly supported perturbation of the constant-coefficient background. In particular, the corresponding Weyl function admits the meromorphic continuation described in Theorem˜10.2.
Corollary 11.1 (Fixed-height Weyl sampling is exponentially ill-conditioned in the full canonical-system class).
Fix $\eta>0$ and sample points $x_{1},\dots,x_{M}\in\mathbb{R}$. For $z=x+i\eta$ write $m(z)$ for a Weyl function and set
$f(x):=\frac{1}{\pi}\operatorname{Im}m(x+i0),\qquad g_{\eta}(x):=\frac{1}{\pi}\operatorname{Im}m(x+i\eta).$
If in Theorem˜10.8 one has $c_{2}=0$ and $d\rho(t)=f(t)\,dt$, then combining Corollary˜10.9 with $P_{\eta}(x)=\frac{1}{\pi}\frac{\eta}{x^{2}+\eta^{2}}$ yields $g_{\eta}=\mathcal{P}_{\eta}f$. Then the map
$m\ \longmapsto\ \bigl(g_{\eta}(x_{k})\bigr)_{k=1}^{M}$
is exponentially ill-conditioned for recovering $f$ in the deterministic minimax sense of Theorem˜10.12: for every $K\geq 2$ and $\varepsilon>0$ there exist Herglotz functions $m^{\pm}$ with boundary densities $f^{\pm}$ (and the constants/functions $c_{1},h_{K}$ from Theorem˜10.12) such that
$\|(g_{\eta}^{+}(x_{k})-g_{\eta}^{-}(x_{k}))_{k=1}^{M}\|_{2}\ \leq\ c_{1}\,\varepsilon\,e^{-\eta(K-1)},\qquad\|f^{+}-f^{-}\|_{L^{2}(\mathbb{R})}\ \geq\ \varepsilon\,\|h_{K}\|_{L^{2}(\mathbb{R})},$
and consequently any reconstruction from noisy fixed-height samples suffers the lower bound of Theorem˜10.12. By the de Branges correspondence Theorem˜10.1, such $m^{\pm}$ are realized as Weyl functions of trace-normed canonical systems.
Proof.
For the two-point alternatives $m^{\pm}$ produced by Theorem˜10.12 (realized via Theorem˜10.1), one has $c_{2}=0$ and $d\rho^{\pm}(t)=f^{\pm}(t)\,dt$, hence Corollary˜10.9 and the definition of $P_{\eta}$ give $g_{\eta}^{\pm}=\mathcal{P}_{\eta}f^{\pm}$. The two-point construction and deterministic minimax lower bound are exactly Theorem˜10.12. Finally, Theorem˜10.1 realizes the constructed Herglotz functions as Weyl functions of trace-normed canonical systems. ∎
Remark 11.2 (Translation to other one-dimensional operators).
Many one-dimensional selfadjoint problems (including Sturm–Liouville/Schrödinger and Dirac systems with scalar potentials) can be realized as canonical systems, and their Weyl functions coincide with canonical-system Weyl functions after standard normalizations; see for instance [6, 1, 18]. For compactly supported perturbations of a free background, one obtains free-tail canonical systems, so the meromorphicity constraint of Remark 10.3 applies. Fixed-height sampling still entails Poisson damping of high-frequency spectral features, but the quantitative minimax construction of Theorem˜10.12 is proved for the full class of trace-normed canonical systems (via the de Branges correspondence).